A Bibliography of Publications in Unix/World, UnixWorld’s Open Computing and Open Computing

Magazines

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

06 September 2011 Version 1.08

Title word cross-reference

1 [Far92a]. 1-2-3 [Far92a]. 100 [KB95]. 1170-based [Tri95w]. 1991 [Poo91d]. 1992 [Won92].

25th [Kir95j].

3 [Far92i, Eng95a]. 3-D [Far92i, Eng95a].

’93 [HD93].

Abandoned [Bal92c]. able [Tri95i]. Access [Hea91]. accessibility [Poo95k]. Accounting [Tri94c]. ACE [Ano92t]. actuated [BM95]. Administering [Bru95d].

Administration [Hay92g]. Adoption [Roh95]. Advanced [Won95e, Eng95g]. Advantage [Ano93w]. After [Hay93]. Against [Poo92a]. Age [RM91]. Agent [Ste95c, Gre95e]. Ahead [Lac92, Poo92g]. air [RF95]. Along [Swa92]. Alpha [Tay93]. already [Tri95e]. Alternative [Ano95e]. always [Ano92-42]. AMD [Lei93]. American [RF95]. amid [KB95]. analysis [Gil95, Gou94b]. analytical [Bae95b]. Andy [Poo92a]. Answers [Ano94a]. Appear [Kir95c]. Application [Lin95b]. Applications [Gra93a, Len95a]. Approach [Mur91, Bur92c]. approaches [RF95]. apps [Gar95a, Gar95a]. Aquitaine [Gre92b]. Arbor [Gil95]. Archie [War92]. Architecture [Tri95l]. areas [Mag95b]. aren’t [Tri95i]. ask [Tri95k]. Assistants [Kol95a]. ATG [Ano95y]. ATGs [Ano95-28]. ATM [Fel95]. Attracting [Mag95d]. audio [Eng95f]. audio- [Eng95f]. Auditing [Far92g]. Auspex [Bal92e]. authoring [Cor95b]. automatic [Str95a, Mul95d]. Automation [RM91]. Automounter [Lei93]. Average [BM95]. aversion [Mul95b]. Avoiding [Poo95a].

Back [Bur92f, Rya94]. Back-up [Bur92a]. balancing [Tri95e, Tri95e]. Bank [Cor95a]. Banking [Bru94a]. Barbarians [Kir95m, Far94c]. Bargain [Bur91b]. barometer [Mul95b]. based [Tri95w]. be [Per93b]. becoming [Eng95c, Tri95a]. Before [Len95d]. Behind [Poo93c]. Benchmarks [Pfi94, Yag94]. Best [Sim95a]. Best-of-Breed [Sim95a]. Beta [Tri95a]. bets [Ano92t]. better [Bae95b, Bal91c]. Between [Far91c]. Beyond [Cam91]. Big [Bal91c]. binds [Tri95l]. Black [Moo92a]. Bleeding [Eng95l]. Block [Kir95l, Poo92e]. Blue [Ano92e]. boast [Bru95b]. Books [Gar95b, Ano95b]. boom [Eng95f]. both [Atk95]. Bottlenecks [Bac92, Hay91c]. bought [Eng95e]. Brain [Poo93c]. Breaking [Far94d]. Breed [Sim95a]. Brewing [Bru94b]. bring [Mag95b]. Brings [Wol93]. Broadcasting [Mag95a]. Brought [Bur92b]. Browser [Atk95]. Bruno [Ano94c]. BSD [Far94e]. BSD/ [Far94e]. Budget [Gil95]. budgeting [Gil95]. Bug [Mul95a, Wal91a]. build [Cor95a]. building [Mag95b, CMQ92]. Builds [Sne92]. Bull [Bur92c]. burgeoning [Bae95a].

Business [Bru95e, Hab95]. Butt [Ano95-31]. Buying [Tri94a]. Bypass [Kir95f].

Cabinet [CR94]. Call [Las95b, Str95a].

Can [Ano94v]. Can’t [Cor94].

capabilities [Eng95a, Eng95f]. Car [BC95a]. Care [Rau95b]. Carol [Twe94]. carrier [CHB95]. CASE [Hay91b]. Catch [SP92]. Catch-up [SP92]. Caught [Kir95h]. CBT [Won94b]. CDE [Won95b]. Cellular [Mag95c]. Center [Mag95d, Str95a]. Center/ Netscape [Mag95d]. centers [Bru95b]. Century [Kir95j]. challenge [Ano92-40]. Champion [Gre92a]. Change [Eng92, Ano94v]. characteristic [KB95]. Chase [Ste95d]. Chasing [Lin95f]. Cheap [Far94e]. Check [Dod94a]. Checking [Kla93]. chips [Ano92e]. Choose [Ano95x]. Choosing [Ano95c]. circumvent [Cam95a]. Civil [Ste95a]. class [Sch95a]. Clearing [Roh95]. Client [Tri94c]. Client-Server [Tri94c]. Clients [Rya94]. closer [Ano91p]. clout [Tri95e]. Co [Mag95a]. Cobol [Tri95n]. Color [Twe95, Bur92b]. Comes [Moo92a]. Comfort [Cra95b]. Commerce [Roh95]. Commercial [Len95c]. commitment [Eng95c]. Common [Poo93b]. communities [RF95]. Companies [Poo91d]. company [Mul95b, Eng95h]. compatibility [Atk95]. Complete [Kir95a]. complex [Bae95b]. Complexity [Sch95a]. complicated [Mul95b]. Compound [Lin95f]. Computer [Ava92b]. Computers [Tri94b]. Computing [FS94]. Conciliator [Poo92i]. Connected [Far92b]. Connection [Ano94t]. Consider [Ano95z]. consolidate [Gil95]. Consuming [Kir95g]. Control [Tri94g, Far94a]. Controlling [Hea91]. convenience [Eng95c].

Conversion [Cra94b]. CORBA [Lin95g]. Corp [Rob95]. corporate [Poo95a]. Cost [Kir95n]. Costs [Ano95z]. Counting [Bur92d]. County [Won95a]. Cover [Ano95-31]. CPU [Bac92]. Cracking [Bis92]. Crash [Poo91a]. create [Len95a]. creates [Ste95c]. Cross [Bur93b]. Cross-Platform [Bur93b]. Crowd [Fry95a]. Cure [Poo92d]. curve [Ver95]. Customers [Tay93, Mag95d]. Cyberspace [Bie95]. Cybertroops [Las95b].

D [Far92i, Eng95a]. Data [Cra94b, BC95b]. Database [Bou95b]. Databases [Bae95b, Tri95t]. Dawn [Won94a]. DCE [Ano95-30, Hir93]. Deal [Tri95x]. Dealer [Poo92h]. dearth [Len95a]. DEC [Nee92]. Decision [Far94a, Ano92-42]. Decision-Support [Far94a]. Decisionmaking [CHB95]. decisions [Tri95b]. Declare [Bal92a]. DECwrite [Bur92e]. defects [Gar95a]. Deletis [Rob95]. Delivering [Stu92]. demand [Ver95]. depends [Cor95a]. desktop [Ano92-54, Tri94e]. desktops [Len95b, MS91]. despite [Len95a]. Details [Eng94d]. develop [Poo95a]. developers [Bae95a]. Developing [Gra93a]. Development [Lin95d, Lac92]. Dial [Gra93b]. Dial-up [Gra93b]. Diane [Eng94b]. Did [Won95a]. difference [Tri95i]. Digital [Kol95a, CR95]. Dilemma [Hay92f]. Directory [War92]. Distributed [Eng95l]. Distributing [BC95b]. distributors [Str95a]. Do [Ano94v, Tri95m]. Document [Lin95f]. Documenting [Gug95a]. Documents [Kri95]. Doing [Fry94]. Dollars [Eng94d]. down [Eng95a, Bal91a]. downsizing [Ano92-48]. Draw [Bur91d]. Dreaming [Kir95n]. DSS [Ste95c].

Dual [Mur91]. Dummies [Won95e]. Dump [Hay92h]. duration [BM95]. Dysfunctional [Eng95h].

E-mail [Bur93b, Cam91]. Easier [Far91d]. Easy [Cla94]. Edge [Eng95l]. Editor [Ano91a]. Editors [Far92d, Ano92f]. Education [Eng95d]. Effortlessly [Far92g]. Electronic [Ano94y]. Elf [Gre92b]. Elusive [Sch95b]. Embrace [Fry95a]. emissions [Cam95a]. Empowers [BC95d]. Emulation [Tri94f]. end [Moo92b]. ends [Bae95a]. enough [Eng95b]. Enterprise [Lin95a]. entire [Atk95]. environment [Len95b]. Epic [Gil95]. Epoch [Gra92]. equipment [Tri95k]. Especially [Ano94d]. Essbase [Gil95]. Establishing [Wol95]. Ethernet [Far92b]. Europe [Gre94a]. European [RF95]. Ever [Eng94a]. Everybody [LaP94]. executives [Rob95]. Expect [Len95b]. Expecting [Ano94e]. expensive [Mul95b]. Expert [Bal92b]. Exploding [Hay92a]. Explosion [FH92]. extent [Tri95k]. External [Kir95a]. Extractors [Bou95b]. Eyes [Poo91b].

Face [Eng92]. Facing [Hay92b]. Factor [Ste95g]. Fail [Bru95a]. Fail-Safe [Bru95a]. Failure [Eng95g]. fall [Ano92-59]. Fangs [Moo92b]. Fannie [Cor95b]. far [Eng95b]. Farther [Kir95c]. Fast [Far94b]. Faster [Moo92a]. Fat [Mul95c]. Fax [Bur93c]. feasibility [Ver95]. features [Str95a]. Fed [Eng95c]. few [Tri95i]. File [CR94]. financial [Gil95]. Find [Eng94c]. Finding [Cra95b]. Finds [Len95d]. Firewall [Cra95g]. Firewall-1 [Cra95g]. Firewalls [Tri95o]. Fix [Ava92a]. Flacks [Ano91h]. flak [Ano91h]. Flat [Lin94a]. Flight [Gug95b]. Foiling [Bis92]. Food [Lin95b]. Force [Lin95e].

Forecast [Poo92c]. Forming [Tri95b]. Forte [Cra95f]. Forums [Ano94y, Ano94d]. Fox [Mag95a]. Frames [Ste95a]. France [Gre92a]. Free [Far91b]. Freeway [Kir95i]. Friendly [MS91]. front [Bae95a]. full [Eng95f].

Gamble [Cra94b]. Game [Ano95-27]. Gap [Hay92b]. General [Poo91a]. generation [Cor95a]. Get [Poo94, Nee91a]. Gets [Bru94c]. Getting [Bal91b]. give [Tri95e]. giving [Sch95a, Hei92]. Global [Kin92]. glue [Ano95h]. going [Eng95b, CR95]. Goodbye [CR94]. Got [Ano95-32]. Grabbag [Ano94w]. Graphical [Hay92c]. Graphics [Swa92, Far92c]. Grapples [Swa92]. Great [Poo92i]. Green [Tri94b]. Grounds [Bru95b]. group [Bae95a]. Groups [Eng95g, PW92]. Groupware [Hay92f, Ano95d]. Growing [Gar95e]. GUI [Lac92, Lin95d]. GUIs [Myn92].

Habit [Far94d]. Hands [Ano93g]. Hands-On [Ano93g]. Harder [Eng94a]. Headed [Poo91a]. Headhunter [Len95d]. Health [Rau95b]. Heat [Gar95f]. Heck [Ano95-30]. Heeeeere [Poo93a]. heist [Eng95b]. Heller [Poo92a]. Help [Gar95b, Eng94c]. helped [Gil95]. Here [Fry94]. heterogeneous [Len95b]. Hewlett [Bru95b]. Hey [Str95c]. Hierarchical [Ste95e]. High [Kir95n, Tri95p]. High-End [Ste95b]. High-Speed [Tri95p]. Homer [Mag95a]. Hooking [Far93a]. Hope [Far92j]. Host [Kol95b]. Hot [Mag95a]. House [Roh95]. HP [Mur91]. Huge [Ano94e]. Hunger [Lin95b].

IBM [Bru95b, Sne92]. identity [Poo95a]. ignore [Mul95a]. I’m [Kir95h]. impacts [Cam95a]. implementation [Mul95b]. Implementing [Tri95y]. implications [RF95]. Improve [Far92h]. Improving [Mui92]. inches [Ano91p]. incidence [Eng95b]. include [Str95a]. increase [Cor95a]. increased [Eng95b]. Independence [Bal92a]. individuals [Tri95b]. industry [KB95, Ano91d]. Influential [Ano95b]. information [Rob95]. insiders [Eng95b]. installed [Tri95k]. installed-and [Tri95k]. intellectual [Eng95b]. interactive [Mag95a]. Internal [Kir95l, Ano95e]. Internet [Wol95, Zin94a]. Interoperability [Bru94a]. Introduction [Ano94g]. Intruders [BC95c]. invest [Str95a]. ISDN [Tri95y]. Island [Swa92]. ISO [Tri95m]. Isolation [Lin94b]. Israeli [Gre94e]. Israeli-Style [Gre94e].

Java [Mag95a]. jewels [KB95]. Job [Ano92-40]. Johnson [Ste95c]. Join [Ano95y]. joint [Ano92-54]. Jump [Poo92g].

Keep [Kir95m]. Keeping [Lin94a].

Key [Kin92]. Kick [Cor94].

Kirzner [Ano95g]. Know [Tri95y].

Lacks [Boo92]. Ladies [Fry94].

Lane [Nee91b]. LANs [Tri95j, Tri95p].

Lap [Far91f]. Laptops [Rau94c]. Large [Twe95]. Last [Poo95c]. late [Won95b]. Latest [Boo92]. launch [Ano92-54]. Leap [Lac92]. Learn [Zin92, Ano95-32]. Learned [Yag94]. Learning [Eng95d]. Leasing [Tri94a]. Least [Wal91b]. led [Ano95h]. Lee [Ano94c]. Lepore [Won94a]. Less [Ano95f]. Let [Ano92p]. Letter [Str95c]. Lexus [Rau94c]. liberalisation [RF95]. Library [Zin95a]. License [Bur91a, Tri95q]. Life [Hay93, Nee91b]. Like [Eng93]. Limits [Wes95]. Lines [BC95c]. link [Str95a]. Linux [Kla93]. List [Ano95b]. Lists [Ano94y]. Live [Yag94]. Living [Bur92b]. Load [Tri95e, Tri95e]. Load-balancing [Tri95e, Tri95e]. Loans [Cor95b]. Logout [Ano94h]. Look [Far92a, Bur92f]. Looking [KB95]. Lot [Ano95-32]. Lotus [Bru95d]. Low [Moo92b]. Low-end [Moo92b]. lower [Tri95j, Far92i].

Mac [Str95c]. machines [Poo95b]. Mae [Cor95b]. Mail [Far91d, Cam91]. Mailing [Ano94y]. Mainframe [Far94d]. Make [Ano95w]. Making [Rya94]. Man [Poo93b]. manage [Len95b]. Management [Tri95v, Bru94d]. Managers [Eng95j]. Managing [Gar95d]. many [Eng95b]. Mapping [Ste95f]. Mark [Pfi94]. market [Ano92-42, Poo92c]. marriage [Eng95c]. Mars [Won95c]. Mastering [Cam92]. matter [Won95b]. Matters [Dia95]. Maximizing [Far92c]. may [Tri95i]. Maybe [Won95b]. McCaw [Mag95c]. McNealy [Ano92-41]. Me [Cra94a]. Meanwhile [Gre94a]. measurement [Poo95k]. measures [Cam95a]. Meet [Ano95h, Zin94a]. meetings [Len95b]. Meets [WP93]. Mega [Tri95x]. Mellowing [Poo91c]. Meltdown [Gar95c]. Men [Ano94d]. Menace [Gar95e]. Mercury [Mag95d]. Microsoft [Tri95l, Tri95r]. Microsystems [Mag95a]. Middle [Mul95b]. Midrange [Tri94a]. migrate [Bru95b]. Migrating [Tri94c]. migration [Bru95b]. Millenium [Mul95a]. Millennium [Mul95a]. Minded [Ano94l]. MIPS [Ano92t]. mirrors [KB95]. MIS [Pfi94]. Miss [Pfi94]. models [Eng95f].

Modem [Far91a]. Modern [Rau95b].

Monitoring [Tri95s].

Monitors [Lin95c, Lin95h].

Month [Ano91i]. Mosaic [Rya94]. Most [Bur92g]. motor [CHB95]. Motorola [Mag95b]. Movable [Eng95m]. multidimensional [Gil95]. Multimedia [Eng95d]. Multiport [Bal91d]. Multiprocessing [Len95a]. multithreaded [Len95a]. Must [Kir95j]. my [Eng95a]. Mystery [Wal91a]. Myths [Hay92a].

National [Rob95]. Need [Tri95m].

Negligence [Eng95i].

Negotiating [Won95e]. Netsite [Mag95d]. Netware [Tri95u]. Network [Bru94d, Far94f]. networks [Eng95b]. Nevada [Won95a]. Never [Fel95, Fel95]. News [Ano91d]. Next [O’C94, Poo91c]. NextStep [Mag95c]. Nifty [Zin94c]. No [Bur91b, Ano91h]. Noorda [O’C94]. Notes [Ano95z, Bru95d]. Nothing [Mul95d]. Nova [Cor95a]. Novell [Ano92-54, Tri94d]. NT [Hay93, Far93b]. Numbers [Mag95b].

O* [Kir95e]. O*U [Kir95m]. Object [Sch95a, Tri95t]. object-oriented [Sch95a, Far92k]. Objective [Hay92d]. Objectivity [Mag95b]. Objectivity/ DB [Mag95b]. Objects [Cor95a, LaP94]. Obsolescence [Ano95-27, Gar95d]. Occur [Gom92]. ODBMS [Cor95a]. Off [BC95c]. off-the-shelf [Ano91p]. offer [Bae95a]. Office [RM91]. often [Tri95a]. Oil [Gre92b]. Okay [Eng95e]. OLAP [Ste95c, Bae95b]. Old [Rau95c]. OLE [Fry95a]. OLTP [Bou95b]. On-Site [Rau94a]. One [Len95d]. Online [Bae95b]. only [Sch95a]. Open [Ano94l]. Openers [Ano94p]. Opening [Per93a].

Opportunities [Cra95e]. orchestrating [Rob95]. oriented [Sch95a, Far92k]. OS/ [Cra95d]. OSF [Hir93]. other [Eng95b]. Outsource [Ano95-29]. Outsourcing [Tri95x]. Overwhelms [Bur93c].

Packages [Kir95a]. Packard [Bru95b]. pain [Bae95a]. pain-free [Bae95a]. Paper [Ste95d]. Parallel [Len95a]. Partner [Ano95c]. Partnerships [Poo95c]. Passing [Gar95a]. Passions [Kir95g]. Password [Bis92]. Patience [Mul95c]. Payback [Ano94e]. Payoff [Sch95b]. PBXs [Str95a]. PC [Ano94t, Bur91e]. PC-Unix [Ano94t]. PCMCIA [Tri94e]. PCs [Eng95a]. [Far94e, Cra95d]. DB [Mag95b]. IP [Len95c, Tri95z]. Netscape [Mag95d]. Sunsoft [Cra95g]. people [Ano92p]. Performance [Tri95s]. Perks [Won95e]. Perl [Wal91b]. Persistence [Far92l]. Personal [Kol95a]. PG [Won95d]. phases [BM95]. picks [Ano92e]. Pink [Poo92b]. plan [Atk95]. planning [Gil95]. Platform [Bur93b]. Play [SP92]. Played [Cra95f]. Plays [Poo92h]. Point [Cra95d]. Polishing [Ste95c]. political [Ver95]. Politics [Eng95e]. Popular [Wol93]. Port [Lin94b]. Power [Bal91a, Eng95k]. presence [Cor95b]. Presentations [Cam92]. previously [Mag95b]. Price [Far92i]. prices [Ano92-59]. pricing [Ver95]. Prime [Bur91c]. Printers [Twe95]. probably [Poo95a]. problem [Mul95a]. Problems [Gom92]. Process [Hab95]. Processes [Bru95e]. processing [Lin95c]. Product [Ano91i]. Productivity [Hay92d]. Products [Bal92d, Ano95w]. Professional [Gar95e]. Profile [Lin94a]. program [Tri95a]. Programmers [Tri95g].

Programming [Man95]. Project [Ano95f]. Projects [Zin94c]. promises [Tri95e]. property [Eng95b]. Proportions [Gra92]. protect [Eng95b]. Protecting [Far94f]. Protection [Ran94]. Protocol [Tri95v]. Proves [Cra95d]. provide [Lin95c]. Proving [Bru95b]. Published [Bor92]. pure [Bae95a]. Pushing [Wes95]. put [Ano92p].

Quality [Gar95a]. queries [Bae95b]. question [Eng95a]. Questions [Bru95a]. Quick [Bur91d].

RAD [Wes95]. radical [Rob95]. RAID [Far92e]. Rather [Far91e]. RDBMS [Cor94]. Re [Eng95m]. reaches [Ano92-42]. Ready [Bru95c]. Ready-to-Use [Bru95c]. Real [Bur92i]. Real-Time [Bur92i]. Reality [Dod94a]. Really [Tri95m]. redesign [Ano95h]. Rediscovers [Tri94d]. Reengineering [Hab95]. regarded [Won95b]. Relating [Kri95]. relational [Bae95b]. relatively [Bae95a]. replace [Tri95i]. Report [Ano91m]. reports [Ste95c, Bru94b]. repositories [Gil95]. Rescue [Far92e]. Reshaping [Bru95e]. Residence [Cra95c]. Resistance [Wal91b]. Resources [Kol95b]. Rest [Far92f]. restrictions [Cam95a]. Results [Wes94b]. Rethinking [Len95a]. Retrain [Tri95g]. Reviews [Rau94a]. revives [Mag95a]. Rid [Hay91a]. Riddle [Mul95b]. Riding [Eng95l]. Right [Ano95c]. Rights [Ste95a]. Rikki [Ano95g]. RISC [Ano92e]. rise [Len95a, Sim95b]. risk [Mul95b]. risky [Tri95a]. Road [Ano94v, Kir95i]. Robbins [Twe94]. Roebuck [Gil95]. role [Ano92-41]. Roommates [Per92b]. routers [Tri95i]. Routing [Tri95i]. Run [Ava92a].

S.C. [Ste95c]. Safe [Cra95g, Bru95a]. Said [Wes94a, Wes94a]. Salaries [Sim95b]. Sales [Ste95c]. Sam [Eng95c]. SATAN [Str95d]. satellite [Mag95b]. satellite-based [Mag95b]. Say [Fel95, CR94]. Saying [Tay93]. says [Gar95a]. Scanners [Bur92j]. Scanning [Bur92j]. Scotia [Cor95a]. Scotty [Cra94a]. Sear [Gil95]. Sears [Gil95]. Season [Kir95k]. Secrets [Ano94d]. Security [Far92g, Mui92]. Seduced [Bal92c]. See [Lin94c]. Seeing [Cra95e]. Segregating [Kol95b]. Semiconductor [Rob95, Ano95h]. separation [Poo95k]. Sequent [Poo92d]. Serlin [Ano91m]. Server [Tri94c, Bru95c]. Servers [Tri94a]. service [Mag95b, Gre95e]. services [Lin95c]. Serving [Bur91e]. Set [Eng95g]. seven [Won95b]. several [Eng95f]. Shareware [Len95c]. Sharing [Kol95b]. shelf [Ano91p]. Shell [Ano95-27]. Shifting [Eng95k]. ship [Eng95f]. shipments [Ano92-59]. Shopping [Bie95]. Short [Ano95b]. Should [Ano95x]. Shouldn’t [Ano95-29]. Show [Eng93]. Shrewd [Zin94c]. signal [BM95]. Silicon [Far92i, Gre94e]. Simple [Tri95v]. single [Tri95j]. Site [Tri95a, Rau94a]. Sleep [Bur92a]. SLIP [Hei92]. Slow [Nee91b, Roh95]. slowly [Sch95a]. small [Bae95a, Twe95]. Smart [Bru94c]. Smarts [Bal92b]. Smigel [Eng94b]. smoke [KB95]. smoke-and-mirrors [KB95]. SMS [Tri95r]. Snapshot [Won93]. Snapshots [Kin93a]. Soft [Poo92f]. Software [Tri94g, Sim95a]. Solbourne [Boo92]. solution [Roh95]. Solutions [Far92m]. Solve [Won94b]. Solves [Ste95d]. Solving [Gom92]. Some [Cra95b]. something [Mul95a]. Sophisticated [Bru94d]. Southern [BC95d]. Space [Gra92]. SPARCstation [Bur92e].

SPARCstations [Far92m].

sparkling [Bru95b]. spatial [Poo95k].

Spec [Tri95w]. Speed [Tri95p]. spend [Len95b]. Spoil [Far91g]. Spot [Poo92f]. Spreadsheets [Bur92i]. Spreadsheets-Really [Bur92i]. SQLWindows [BC95d]. Squashing [Mul95a]. Squeeze [Hay92g]. Stacks [Len95c]. stake [Cor95b]. Stalking [Tri95x]. standard [Str95a]. Standardize [Tri95u]. standardizing [Tri95j]. Standards [Lin95f]. Star [Eng95h, Ano95w]. Stars [Rau95d]. starters [Atk95]. Stays [Poo92g]. Step [Eng95d]. Steve [Ano92-40]. Stick [Ano95h]. Still [Tri95n, Lei92]. Storage [Bru94c, Gra92]. Storefront [Wol95]. Storm [Bal91d]. Strategy [Sne92]. Street [SWG92]. Stress [Eng95j]. study [Gar95a]. Style [Gre94e]. Success [Far91g, Gug95a]. Successful [Ano95-28]. Suddenly [Eng95c]. suite [Atk95]. Suites [Sim95a]. Sun [Far91e, Poo91b]. supercomputer [Tri95e]. Support [Far94a]. supposed [Tri95l]. Sure [Bal91c]. Survey [Pfi94]. Survives [Ran95]. SVR4 [Hea91]. Switch [Far94b]. Switches [Tri95i]. Switching [Tri95i]. Sybase [Poo92g]. System [Far92h, She92]. Systems [Poo92c, Bru95b].

T [Kir95m]. Take [Str95c, Bur92k]. Takes [Mag95c]. Taking [Gug95b]. Talk [Rya94, BC95a]. Tape [Bal91b]. Target [Ano95f]. TCP [Len95c, Tri95z]. TCP/IP [Len95c, Tri95z]. team [Ano95h]. Techniques [Rau95c]. Technology [Eng95g, Ano92p]. Telecommunications [Cra95f]. telephone [Mag95b]. tell [Tri95i]. Ten [Tri95y]. Test [Gar95a]. Text [Sch95a]. Text-only [Sch95a]. Their [Poo91b].

Them [Poo95c]. themselves [Eng95b]. There [Hay93]. Thin [Mul95c]. Things [Tri95y]. Think [Mul95a]. Thoughts [Ano91a]. Threat [Kir95l]. Three [Fry95b]. Three-Tier [Fry95b]. tie [Tri95l]. Tier [Fry95b]. Time [Kir95h, Cra95g]. Time/ Sunsoft [Cra95g]. Tires [Cor94]. Tofu [Poo95a]. Toll [Kir95i]. Tommy [Lin94c]. too [Won95b]. Tool [Far94a]. Tools [Hay92e, Hab95]. Top [Poo94]. Toxic [BC95b]. Toy [Bur91b]. TP [Lin95h]. Track [Far94b]. trade [Eng95b]. Training [Won94b, Won95d]. Transaction [Lin95c]. Transform [Wol93]. Transition [Nee92]. Translating [Myn92]. transport [Tri95j]. transportation [Poo95k]. Tries [Bur92c]. Trouble [Tri95z]. truck [Cam95a]. trucks [Cam95a]. Trust [Eng95e]. trustworthy [Lin95c]. TSAPI [Tri95k]. Turn [Far94g]. Turning [Gar95f]. Turnkey [Bur92c]. Tutorial [Lei93]. twice [Gar95a]. Two [SWG92, Far92m]. Two-Way [SWG92]. types [Cor95a].

U [Kir95e]. U.S. [Gre93]. Unbundled [Hay92e]. Uncle [Eng95c]. Uncovering [Bac92]. Understanding [Far93b]. Uneasy [Far94e]. UniSQL [Ste95d]. Unix [Tri94d, Ano91p]. UnixWorld [Hay92h]. Unjamming [Hay91c]. Upgrade [Lin95e]. Upheaval [Rob95]. upside [Ano92-48]. Use [Bru95c]. User [PW92]. Users [Won95c]. Using [Lei93, Hir93]. USL [Ano92-54]. Utilities [Zin94c].

Valley [Gre94e]. Vary [Wes94b]. Vendor [Gom92]. Vendors [Won95c]. venture [Tri95a, Ano92-54]. ventured [Mag95b]. Venus [Won95c]. Versant [Cor95a]. Version [Tri94g].

Video [Far92m].

video-processing [Eng95f].

videoconferencing [Eng95f, Eng95f].

vision [Ano92-40, Far92l].

Visionary [Man95]. Visualization [Wol93]. Volume [Far94g]. vs [Tri94f].

Wait [Kir95j]. Waiting [Ano95a]. Walls [Far94c]. Wanna [Zin92]. WANs [Kin92]. Want [Ano95y]. Warp [Cra95d]. was [Roh95]. Wax [Ste95c]. Way [Bur91a]. Web [Ano95e, Cra95b]. WebForce [Cor95b]. well [Won95b]. well-regarded [Won95b]. Where [Eng94c, Sch95b]. whether [Tri95k]. which [Tri95k, Bal91d]. While [Bur92a]. Who [Tri95b]. Wide [Poo95a]. Will [Wes94b, Won94b]. win [Ste95c]. Windfall [Lei92]. Window [She92]. Windows [Tri94f, Tri94f]. Wireless [Bru95f]. Wires [Hay91a]. Wizard [Bal92e, Ano94w]. Woes [Won94b]. Women [FS94, Ano94y]. Wordperfect [Eng92]. Work [Hay91b]. worker [Cor95a, Eng95m]. Works [Bur92a]. Workstation [Far91f, Ano92-59]. Workstations [Moo92b, Mur91]. World [Poo92a]. Worldblazer [Gra93b]. Worldwide [Gra93b]. Worth [Ano95a]. WOSA [Tri95l]. Would [Far91e]. Wrong [Mag95b].

X [Ste95g, Per92b]. X.25 [Kin92].

year [Eng95f]. years [Won95b]. Yer [Far91b]. yetbut [Tri95i]. you’ve [Eng95e, Ano95-32].

References

Anonymous:1991:ETa

[Ano91a] Anonymous. Editor’s Thoughts. UNIX/world, 8(10):11–??, October 1991. ISSN 0739-5922.

Anonymous:1991:ETb

[Ano91b] Anonymous. Editor’s Thoughts.

UNIX/world, 8(11):11–??, November 1991. ISSN 0739-5922.

Anonymous:1991:ETc

[Ano91c] Anonymous. Editor’s Thoughts.

UNIX/world, 8(12):11–??, December 1991. ISSN 0739-5922.

Anonymous:1991:INd

[Ano91d] Anonymous. Industry News.

UNIX/world, 8(9):13–??, September 1991. ISSN 0739-5922.

Anonymous:1991:INa

[Ano91e] Anonymous. Industry News.

UNIX/world, 8(10):15–??, October 1991. ISSN 0739-5922.

Anonymous:1991:INb

[Ano91f] Anonymous. Industry News.

UNIX/world, 8(11):15–??, November 1991. ISSN 0739-5922.

Anonymous:1991:INc

[Ano91g] Anonymous. Industry News.

UNIX/world, 8(12):15–??, December 1991. ISSN 0739-5922.

Anonymous:1991:NMF

[Ano91h] Anonymous. No more flak for Flacks. UNIX/world, 8(9):9–??, September 1991. ISSN 0739-5922.

Anonymous:1991:PMd

[Ano91i] Anonymous. Product of the Month. UNIX/world, 8(9):41–??, September 1991. ISSN 0739-5922.

Anonymous:1991:PMa

[Ano91j] Anonymous. Product of the Month. UNIX/world, 8(10):41–??, October 1991. ISSN 0739-5922.

Anonymous:1991:PMb

[Ano91k] Anonymous. Product of the Month. UNIX/world, 8(11):45–??, November 1991. ISSN 0739-5922.

Anonymous:1991:PMc

[Ano91l] Anonymous. Product of the Month. UNIX/world, 8(12):41–??, December 1991. ISSN 0739-5922.

Anonymous:1991:SRa

[Ano91m] Anonymous. The Serlin Report. UNIX/world, 8(10):39–??, October 1991. ISSN 0739-5922.

Anonymous:1991:SRb

[Ano91n] Anonymous. The Serlin Report. UNIX/world, 8(11):39–??, November 1991. ISSN 0739-5922.

Anonymous:1991:SRc

[Ano91o] Anonymous. The Serlin Report. UNIX/world, 8(12):35–??, December 1991. ISSN 0739-5922.

Anonymous:1991:UIC

[Ano91p] Anonymous. UNIX inches closer to off-the-shelf software. UNIX/world, 8(9):33–??, September 1991. ISSN 0739-5922.

Anonymous:1992:AUc

[Ano92a] Anonymous. Answers to UNIX. UNIX/world, 9(8):109–??, August 1992. ISSN 0739-5922.

Anonymous:1992:AUd

[Ano92b] Anonymous. Answers to UNIX.

UNIX/world, 9(9):121–??, September 1992. ISSN 0739-5922.

Anonymous:1992:AUa

[Ano92c] Anonymous. Answers to UNIX.

UNIX/world, 9(10):117–??, October 1992. ISSN 0739-5922.

Anonymous:1992:AUb

[Ano92d] Anonymous. Answers to UNIX.

UNIX/world, 9(11):127–??, November 1992. ISSN 0739-5922.

Anonymous:1992:BPB

[Ano92e] Anonymous. Bull picks Big Blue for RISC chips. UNIX/world, 9 (4):13–??, April 1992. ISSN 0739-5922.

Anonymous:1992:ETe

[Ano92f] Anonymous. Editors Thoughts.

UNIX/world, IX(12):11–??, December 1992. ISSN 0739-5922.

Anonymous:1992:ETc

[Ano92g] Anonymous. Editor’s Thoughts.

UNIX/world, 9(8):9–??, August 1992. ISSN 0739-5922.

Anonymous:1992:ETd

[Ano92h] Anonymous. Editor’s Thoughts.

UNIX/world, 9(9):9–??, September 1992. ISSN 0739-5922.

Anonymous:1992:ETa

[Ano92i] Anonymous. Editor’s Thoughts.

UNIX/world, 9(10):11–??, October 1992. ISSN 0739-5922.

Anonymous:1992:ETb

[Ano92j] Anonymous. Editor’s Thoughts. UNIX/world, 9(11):11–??, November 1992. ISSN 0739-5922.

Anonymous:1992:INe

[Ano92k] Anonymous. Industry News. UNIX/world, IX(12):15–??, December 1992. ISSN 0739-5922.

Anonymous:1992:INc

[Ano92l] Anonymous. Industry News. UNIX/world, 9(8):13–??, August 1992. ISSN 0739-5922.

Anonymous:1992:INd

[Ano92m] Anonymous. Industry News. UNIX/world, 9(9):13–??, September 1992. ISSN 0739-5922.

Anonymous:1992:INa

[Ano92n] Anonymous. Industry News. UNIX/world, 9(10):15–??, October 1992. ISSN 0739-5922.

Anonymous:1992:INb

[Ano92o] Anonymous. Industry News. UNIX/world, 9(11):15–??, November 1992. ISSN 0739-5922.

Anonymous:1992:LPP

[Ano92p] Anonymous. Let’s put people ahead of technology. UNIX/world, 9(4):9–??, April 1992. ISSN 0739-5922.

Anonymous:1992:Lb

[Ano92q] Anonymous. Library. UNIX/world, 9(8):27–??, August 1992. ISSN 0739-5922.

Anonymous:1992:Lc

[Ano92r] Anonymous. Library. UNIX/world, 9(9):27–??, September 1992. ISSN 0739-5922.

Anonymous:1992:La

[Ano92s] Anonymous. Library. UNIX/world, 9(11):31–??, November 1992. ISSN 0739-5922.

Anonymous:1992:MBA

[Ano92t] Anonymous. MIPS bets on ACE.

UNIX/world, 9(1):13–??, January 1992. ISSN 0739-5922.

Anonymous:1992:NPf

[Ano92u] Anonymous. New Products.

UNIX/world, IX(12):139–??, December 1992. ISSN 0739-5922.

Anonymous:1992:NPc

[Ano92v] Anonymous. New Products.

UNIX/world, 9(7):117–??, July 1992. ISSN 0739-5922.

Anonymous:1992:NPd

[Ano92w] Anonymous. New Products.

UNIX/world, 9(8):115–??, August 1992. ISSN 0739-5922.

Anonymous:1992:NPe

[Ano92x] Anonymous. New Products.

UNIX/world, 9(9):127–??, September 1992. ISSN 0739-5922.

Anonymous:1992:NPa

[Ano92y] Anonymous. New Products.

UNIX/world, 9(10):125–??, October 1992. ISSN 0739-5922.

Anonymous:1992:NPb

[Ano92z] Anonymous. New Products.

UNIX/world, 9(11):135–??, November 1992. ISSN 0739-5922.

Anonymous:1992:PMa

[Ano92-27] Anonymous. Product of the Month. UNIX/world, 9(1):41–??, January 1992. ISSN 0739-5922.

Anonymous:1992:PMc

[Ano92-28] Anonymous. Product of the Month. UNIX/world, 9(2):35–??, February 1992. ISSN 0739-5922.

Anonymous:1992:PMd

[Ano92-29] Anonymous. Product of the Month. UNIX/world, 9(3):31–??, March 1992. ISSN 0739-5922.

Anonymous:1992:PMe

[Ano92-30] Anonymous. Product of the Month. UNIX/world, 9(4):35–??, April 1992. ISSN 0739-5922.

Anonymous:1992:PMf

[Ano92-31] Anonymous. Product of the Month. UNIX/world, 9(7):41–??, July 1992. ISSN 0739-5922.

Anonymous:1992:PMg

[Ano92-32] Anonymous. Product of the Month. UNIX/world, 9(8):43–??, August 1992. ISSN 0739-5922.

Anonymous:1992:PMh

[Ano92-33] Anonymous. Product of the Month. UNIX/world, 9(9):41–??, September 1992. ISSN 0739-5922.

Anonymous:1992:PMb

[Ano92-34] Anonymous. Product of the Month. UNIX/world, 9(10):43–??, October 1992. ISSN 0739-5922.

Anonymous:1992:Se

[Ano92-35] Anonymous. Standards. UNIX/world, IX(12):111–??, December 1992. ISSN 0739-5922.

Anonymous:1992:Sc

[Ano92-36] Anonymous. Standards. UNIX/world, 9(8):89–??, August 1992. ISSN 0739-5922.

Anonymous:1992:Sd

[Ano92-37] Anonymous. Standards. UNIX/world, 9(9):95–??, September 1992. ISSN 0739-5922.

Anonymous:1992:Sa

[Ano92-38] Anonymous. Standards. UNIX/world, 9(10):97–??, October 1992. ISSN 0739-5922.

Anonymous:1992:Sb

[Ano92-39] Anonymous. Standards. UNIX/world, 9(11):99–??, November 1992. ISSN 0739-5922.

Anonymous:1992:SJV

[Ano92-40] Anonymous. Steve Jobs’ vision and challenge. UNIX/world, 9(1): 9–??, January 1992. ISSN 0739-5922.

Anonymous:1992:SSM

[Ano92-41] Anonymous. Sun’s Scott McNealy on a role. UNIX/world, 9(2):39– ??, February 1992. ISSN 0739-5922.

Anonymous:1992:MAR

[Ano92-42] Anonymous. The market always reaches a decision. UNIX/world, 9 (3):9–??, March 1992. ISSN 0739-

Anonymous:1992:SRe

[Ano92-43] Anonymous. The Serlin Report. UNIX/world, IX(12):37–??, December 1992. ISSN 0739-5922.

Anonymous:1992:SRc

[Ano92-44] Anonymous. The Serlin Report. UNIX/world, 9(8):39–??, August 1992. ISSN 0739-5922.

Anonymous:1992:SRd

[Ano92-45] Anonymous. The Serlin Report. UNIX/world, 9(9):37–??, September 1992. ISSN 0739-5922.

Anonymous:1992:SRa

[Ano92-46] Anonymous. The Serlin Report. UNIX/world, 9(10):39–??, October 1992. ISSN 0739-5922.

Anonymous:1992:SRb

[Ano92-47] Anonymous. The Serlin Report. UNIX/world, 9(11):41–??, November 1992. ISSN 0739-5922.

Anonymous:1992:UD

[Ano92-48] Anonymous. The upside to downsizing. UNIX/world, 9(2):9–??, February 1992. ISSN 0739-5922.

Anonymous:1992:UEe

[Ano92-49] Anonymous. UNIX in Europe. UNIX/world, IX(12):31–??, December 1992. ISSN 0739-5922.

Anonymous:1992:UEc

[Ano92-50] Anonymous. UNIX in Europe. UNIX/world, 9(8):33–??, August 1992. ISSN 0739-5922.

Anonymous:1992:UEd

[Ano92-51] Anonymous. UNIX in Europe. UNIX/world, 9(9):31–??, September 1992. ISSN 0739-5922.

Anonymous:1992:UEa

[Ano92-52] Anonymous. UNIX in Europe.

UNIX/world, 9(10):31–??, October 1992. ISSN 0739-5922.

Anonymous:1992:UEb

[Ano92-53] Anonymous. UNIX in Europe. UNIX/world, 9(11):35–??, November 1992. ISSN 0739-5922.

Anonymous:1992:UNL

[Ano92-54] Anonymous. USL and Novell launch joint desktop software venture. UNIX/world, 9(3):13–??, March 1992. ISSN 0739-5922.

Anonymous:1992:WGc

[Ano92-55] Anonymous. Wizard’s Grabbag.

UNIX/world, 9(8):101–??, August 1992. ISSN 0739-5922.

Anonymous:1992:WGd

[Ano92-56] Anonymous. Wizard’s Grabbag. UNIX/world, 9(9):107–??, September 1992. ISSN 0739-5922.

Anonymous:1992:WGa

[Ano92-57] Anonymous. Wizard’s Grabbag.

UNIX/world, 9(10):107–??, October 1992. ISSN 0739-5922.

Anonymous:1992:WGb

[Ano92-58] Anonymous. Wizard’s Grabbag. UNIX/world, 9(11):117–??, November 1992. ISSN 0739-5922.

Anonymous:1992:WSF

[Ano92-59] Anonymous. Workstation shipments fall, but so do prices. UNIX/world, 9(2):13–??, February 1992. ISSN 0739-5922.

Anonymous:1993:AUa

[Ano93a] Anonymous. Answers to UNIX. UNIX/world, 10(1):91–??, January 1993. ISSN 0739-5922.

Anonymous:1993:AUb

[Ano93b] Anonymous. Answers to UNIX. UNIX/world, 10(2):79–??, February 1993. ISSN 0739-5922.

Anonymous:1993:AUc

[Ano93c] Anonymous. Answers to UNIX. UNIX/world, 10(3):87–??, March 1993. ISSN 0739-5922.

Anonymous:1993:ETa

[Ano93d] Anonymous. Editor’s Thoughts. UNIX/world, 10(1):11–??, January 1993. ISSN 0739-5922.

Anonymous:1993:ETb

[Ano93e] Anonymous. Editor’s Thoughts. UNIX/world, 10(2):11–??, February 1993. ISSN 0739-5922.

Anonymous:1993:ETc

[Ano93f] Anonymous. Editor’s Thoughts. UNIX/world, 10(3):11–??, March 1993. ISSN 0739-5922.

Anonymous:1993:HIa

[Ano93g] Anonymous. Hands-On Introduction. UNIX/world, 10(1):70–??, January 1993. ISSN 0739-5922.

Anonymous:1993:HIb

[Ano93h] Anonymous. Hands-On Introduction. UNIX/world, 10(2):62–??, February 1993. ISSN 0739-5922.

Anonymous:1993:HIc

[Ano93i] Anonymous. Hands-On Introduction. UNIX/world, 10(3):66–??, March 1993. ISSN 0739-5922.

Anonymous:1993:INa

[Ano93j] Anonymous. Industry News.

UNIX/world, 10(1):15–??, January 1993. ISSN 0739-5922.

Anonymous:1993:INb

[Ano93k] Anonymous. Industry News.

UNIX/world, 10(2):15–??, February 1993. ISSN 0739-5922.

Anonymous:1993:INc

[Ano93l] Anonymous. Industry News.

UNIX/world, 10(3):15–??, March 1993. ISSN 0739-5922.

Anonymous:1993:Lc

[Ano93m] Anonymous. Library. UNIX/world, 10(2):33–??, February 1993. ISSN 0739-5922.

Anonymous:1993:Le

[Ano93n] Anonymous. Library. UNIX/world, 10(3):33–??, March 1993. ISSN 0739-5922.

Anonymous:1993:La

[Ano93o] Anonymous. Logout. UNIX/world, 10(1):128–??, January 1993. ISSN 0739-5922.

Anonymous:1993:Lb

[Ano93p] Anonymous. Logout. UNIX/world, 10(2):126–??, February 1993. ISSN 0739-5922.

Anonymous:1993:Ld

[Ano93q] Anonymous. Logout. UNIX/world, 10(3):136–??, March 1993. ISSN 0739-5922.

Anonymous:1993:NP

[Ano93r] Anonymous. New Products. UNIX/world, 10(1):115–??, January 1993. ISSN 0739-5922.

Anonymous:1993:Sa

[Ano93s] Anonymous. Standards. UNIX/world, 10(1):39–??, January 1993. ISSN 0739-5922.

Anonymous:1993:Sb

[Ano93t] Anonymous. Standards. UNIX/world, 10(2):29–??, February 1993. ISSN 0739-5922.

Anonymous:1993:Sc

[Ano93u] Anonymous. Standards. UNIX/world, 10(3):29–??, March 1993. ISSN 0739-5922.

Anonymous:1993:SR

[Ano93v] Anonymous. The Serlin Report. UNIX/world, 10(1):31–??, January 1993. ISSN 0739-5922.

Anonymous:1993:XAa

[Ano93w] Anonymous. The X Advantage. UNIX/world, 10(1):97–??, January 1993. ISSN 0739-5922.

Anonymous:1993:XAb

[Ano93x] Anonymous. The X Advantage. UNIX/world, 10(2):85–??, February 1993. ISSN 0739-5922.

Anonymous:1993:XAc

[Ano93y] Anonymous. The X Advantage. UNIX/world, 10(3):91–??, March 1993. ISSN 0739-5922.

Anonymous:1993:UEa

[Ano93z] Anonymous. UNIX in Europe. UNIX/world, 10(1):39–??, January 1993. ISSN 0739-5922.

Anonymous:1993:UEb

[Ano93-27] Anonymous. UNIX in Europe.

UNIX/world, 10(2):37–??, February 1993. ISSN 0739-5922.

Anonymous:1993:UEc

[Ano93-28] Anonymous. UNIX in Europe.

UNIX/world, 10(3):37–??, March 1993. ISSN 0739-5922.

Anonymous:1993:WGa

[Ano93-29] Anonymous. Wizard’s Grabbag.

UNIX/world, 10(1):83–??, January 1993. ISSN 0739-5922.

Anonymous:1993:WGb

[Ano93-30] Anonymous. Wizard’s Grabbag.

UNIX/world, 10(2):75–??, February 1993. ISSN 0739-5922.

Anonymous:1993:WGc

[Ano93-31] Anonymous. Wizard’s Grabbag.

UNIX/world, 10(3):77–??, March 1993. ISSN 0739-5922.

Anonymous:1994:AUa

[Ano94a] Anonymous. Answers to Unix.

Open Computing, 11(10):110–??, October 1994. CODEN OP-COEB. ISSN 1078-2370.

Anonymous:1994:AUb

[Ano94b] Anonymous. Answers to Unix.

UnixWorld’s Open Computing, 11 (9):109–??, September 1994. CO-DEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:BL

[Ano94c] Anonymous. Bruno, Lee. Unix-World’s Open Computing, 11(9): 63–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:EMS

[Ano94d] Anonymous. Especially for Men: Secrets of Women’s Forums. Open Computing, 11(12):63–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:EHP

[Ano94e] Anonymous. Expecting a Huge Payback. Open Computing, 11 (10):45–??, October 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:HHI

[Ano94f] Anonymous. Hands-on help: Introduction. UnixWorld’s Open Computing, 11(9):93–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:I

[Ano94g] Anonymous. Introduction. Open Computing, 11(10):97–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:La

[Ano94h] Anonymous. Logout. Open Computing, 11(10):120–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:Lb

[Ano94i] Anonymous. Logout. Open Computing, 11(11):120–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:Lc

[Ano94j] Anonymous. Logout. Open Computing, 11(12):112–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:Ld

[Ano94k] Anonymous. Logout. Unix-World’s Open Computing, 11(9): 120–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:OMa

[Ano94l] Anonymous. Open Minded. Open Computing, 11(10):11–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:OMb

[Ano94m] Anonymous. Open Minded. Open Computing, 11(11):9–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:OMc

[Ano94n] Anonymous. Open Minded. Open Computing, 11(12):9–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:OMd

[Ano94o] Anonymous. Open Minded. Unix-World’s Open Computing, 11(9): 11–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:Oa

[Ano94p] Anonymous. Openers. Open Computing, 11(10):21–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:Ob

[Ano94q] Anonymous. Openers. Open Computing, 11(11):21–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:Oc

[Ano94r] Anonymous. Openers. Open Computing, 11(12):19–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:Od

[Ano94s] Anonymous. Openers. Unix-World’s Open Computing, 11(9): 21–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:PCa

[Ano94t] Anonymous. The PC-Unix Connection. Open Computing, 11(10): 113–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:PCb

[Ano94u] Anonymous. The PC-Unix Connection. UnixWorld’s Open Computing, 11(9):113–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:WMC

[Ano94v] Anonymous. What Men Can Do: The Road to Change. Open Computing, 11(12):63–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1994:WGa

[Ano94w] Anonymous. Wizard’s Grabbag. Open Computing, 11(10):105–??, October 1994. CODEN OP-COEB. ISSN 1078-2370.

Anonymous:1994:WGb

[Ano94x] Anonymous. Wizard’s Grabbag. UnixWorld’s Open Computing, 11 (9):101–??, September 1994. CO-DEN OPCOEB. ISSN 1072-4044.

Anonymous:1994:WEF

[Ano94y] Anonymous. Women’s Electronic Forums and Mailing Lists. Open Computing, 11(12):66–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:DWW

[Ano95a] Anonymous. A DCE Worth Waiting For. Open Computing, 12(2): 43–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:SLI

[Ano95b] Anonymous. A Short List of Influential Management Books. Open Computing, 12(6):46–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:CRP

[Ano95c] Anonymous. Choosing the Right Partner. Open Computing, 12(3): 39–??, March 1995. CODEN OP-COEB. ISSN 1078-2370.

Anonymous:1995:GWI

[Ano95d] Anonymous. Groupware: What It Is. UnixWorld’s Open Computing, 12(9):32–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Anonymous:1995:IWN

[Ano95e] Anonymous. Internal Web: A Notes Alternative. UnixWorld’s Open Computing, 12(9):40–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1995:KYP

[Ano95f] Anonymous. Keeping Your Project (More or Less) on Target. Open Computing, 12(1):43– ??, January 1995. CODEN OP-COEB. ISSN 1078-2370.

Anonymous:1995:KR

[Ano95g] Anonymous. Kirzner, Rikki. Open Computing, 12(3):79–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:MTS

[Ano95h] Anonymous. Making Things Stick: Meet the glue team that has led the IS redesign at National Semiconductor. Unix-World’s Open Computing, 12(12): 45–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1995:OMa

[Ano95i] Anonymous. Open Minded. Open Computing, 12(1):9–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:OMb

[Ano95j] Anonymous. Open Minded. Open Computing, 12(2):9–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:OMc

[Ano95k] Anonymous. Open Minded. Open Computing, 12(3):9–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:OMd

[Ano95l] Anonymous. Open Minded. Open Computing, 12(4):9–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:OMe

[Ano95m] Anonymous. Open Minded. Open Computing, 12(5):9–??, May 1995.

CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:OMf

[Ano95n] Anonymous. Open Minded. Open Computing, 12(6):9–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:OMg

[Ano95o] Anonymous. Open Minded. Open Computing, 12(7):9–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:Oa

[Ano95p] Anonymous. Openers. Open Computing, 12(1):15–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:Ob

[Ano95q] Anonymous. Openers. Open Computing, 12(2):15–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:Oc

[Ano95r] Anonymous. Openers. Open Computing, 12(3):15–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:Od

[Ano95s] Anonymous. Openers. Open Computing, 12(4):15–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:Oe

[Ano95t] Anonymous. Openers. Open Computing, 12(5):19–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:Of

[Ano95u] Anonymous. Openers. Open Computing, 12(6):15–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:Og

[Ano95v] Anonymous. Openers. Open Computing, 12(7):15–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:PCM

[Ano95w] Anonymous. Products That Can Make You a Star. Open Computing, 12(5):35–??, May 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:SYC

[Ano95x] Anonymous. Should You Choose Outsourcing? Open Computing, 12(3):34–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:YWJ

[Ano95y] Anonymous. So You Want to Join an ATG. Open Computing, 12(6): 42–??, June 1995. CODEN OP-COEB. ISSN 1078-2370.

Anonymous:1995:TNC

[Ano95z] Anonymous. Taking Notes? Consider the Costs. UnixWorld’s Open Computing, 12(9):36–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Anonymous:1995:OSG

[Ano95-27] Anonymous. The Obsolescence Shell Game. Open Computing, 12 (3):47–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:WDS

[Ano95-28] Anonymous. What Do Successful ATGs Do? Open Computing, 12(6):41–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:WSY

[Ano95-29] Anonymous. What Shouldn’t You Outsource? Open Computing, 12 (3):40–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:WHI

[Ano95-30] Anonymous. What the Heck Is DCE? Open Computing, 12(2):39– ??, February 1995. CODEN OP-COEB. ISSN 1078-2370.

Anonymous:1995:WDW

[Ano95-31] Anonymous. What to Do, What Not to Do… And When: How to Cover Your Butt. Open Computing, 12(1):40–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Anonymous:1995:YGL

[Ano95-32] Anonymous. You’ve Got a Lot to Learn. Open Computing, 12(2): 37–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Atkins:1995:IWB

[Atk95] S. Atkins. An Internet Web Browser: For starters, think about compatibility in the entire suite of tools you plan to use, both on the World Wide Web and the Internet. UnixWorld’s Open Computing, 12(11):51–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Avant:1992:RFU

[Ava92a] Harry Avant. Run DOS to Fix UNIX. UNIX/world, 9(10):95–??, October 1992. ISSN 0739-5922.

Avant:1992:WTY

[Ava92b] Harry Avant. Why Not Talk to Your Computer? UNIX/world, 9 (9):81–??, September 1992. ISSN 0739-5922.

Backman:1992:UCB

[Bac92] Rex Backman. Uncovering CPU Bottlenecks with UNIX Tools. UNIX/world, 9(10):101–??, October 1992. ISSN 0739-5922.

Baer:1995:ODD

[Bae95a] T. Baer. Object-Oriented Database Development Tools: A small but burgeoning group of tools developers offer “pure” and relatively pain-free object front ends. UnixWorld’s Open Computing, 12 (10):54–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Baer:1995:ODO

[Bae95b] T. Baer. OLAP Databases?: Online analytical processing is proving to be better at complex queries than relational tools. Unix-World’s Open Computing, 12(9): 60–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Baldwin:1991:GDW

[Bal91a] Howard Baldwin. Getting Data Up When Power is Down. UNIX/world, 8(10):91–??, October 1991. ISSN 0739-5922.

Baldwin:1991:GIT

[Bal91b] Howard Baldwin. Getting It On Tape. UNIX/world, 8(9):91–??, September 1991. ISSN 0739-5922.

Baldwin:1991:MSB

[Bal91c] Howard Baldwin. Making Sure Big Really is Better. UNIX/world, 8(12):69–??, December 1991. ISSN 0739-5922.

Baldwin:1991:WMS

[Bal91d] Howard Baldwin. Which Multiport in the Storm? UNIX/world, 8(11):83–??, November 1991. ISSN 0739-5922.

Baldwin:1992:DYD

[Bal92a] Howard Baldwin. Declare Your Database Independence. UNIX/world, 9(3):59–??, March 1992. ISSN 0739-5922.

Baldwin:1992:ESC

[Bal92b] Howard Baldwin. Expert Systems: Computing with Smarts. UNIX/world, 9(3):47–??, March 1992. ISSN 0739-5922.

Baldwin:1992:SA

[Bal92c] Howard Baldwin. Seduced and Abandoned. UNIX/world, 9(8): 62–??, August 1992. ISSN 0739-5922.

Baldwin:1992:BP

[Bal92d] Howard Baldwin. The Best Products of 1991. UNIX/world, 9(1): 55–??, January 1992. ISSN 0739-5922.

Baldwin:1992:WA

[Bal92e] Howard Baldwin. The Wizard of Auspex. UNIX/world, 9(7):61–??, July 1992. ISSN 0739-5922.

Bruno:1995:CT

[BC95a] Lee Bruno and Gael Core. Car Talk. Open Computing, 12(3):94– ??, March 1995. CODEN OP-COEB. ISSN 1078-2370.

Bruno:1995:DTD

[BC95b] Lee Bruno and Gael Core. Distributing Toxic Data. Open Computing, 12(4):88–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1995:KIL

[BC95c] Lee Bruno and Gael Core. Keeping Intruders Off the Lines. Open Computing, 12(5):88–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1995:SES

[BC95d] Lee Bruno and Gael Core. SQL-Windows Empowers Southern Company. Open Computing, 12 (6):102–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Biesada:1995:SC

[Bie95] Alexandra M. Biesada. Shopping in Cyberspace. Open Computing, 12(2):70–??, February 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Bishop:1992:FPC

[Bis92] Matt Bishop. Foiling Password Cracking. UNIX/world, 9(3):85– ??, March 1992. ISSN 0739-5922.

Bonneson:1995:ADP

[BM95] J. A. Bonneson and P. T. Mc-Coy. Average duration and performance of actuated signal phases. UnixWorld’s Open Computing, 12 (12):429–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Boozer:1992:SLL

[Boo92] Cimarron Boozer. Solbourne’s Latest Lacks Sun Speed. UNIX/world, 9(1):81–??, January 1992. ISSN 0739-5922.

Borsook:1992:YWY

[Bor92]  Paulina Borsook.  So, You

Want Your Software Published? UNIX/world, 9(11):71–??, November 1992. ISSN 0739-5922.

Bournellis:1995:ODEb

[Bou95a] C. Bournellis. OLTP Database Extractors. UnixWorld’s Open Computing, 12(7):53–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Bournellis:1995:ODEa

[Bou95b] Cynthia Bournellis. OLTP Database Extractors. Open Computing, 12(7):53–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1994:BI

[Bru94a] Lee Bruno. Banking on Interoperability. Open Computing, 11(11): 54–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1994:BR

[Bru94b] Lee Bruno. Brewing Reports.

Open Computing, 11(11):109–??, November 1994. CODEN OP-COEB. ISSN 1078-2370.

Bruno:1994:NSG

[Bru94c] Lee Bruno. Network Storage Gets Smart. Open Computing, 11(12): 69–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1994:SNM

[Bru94d] Lee Bruno. Sophisticated Network Management. Open Computing, 11(12):100–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1995:FQA

[Bru95a] L. Bruno. Fail-Safe Questions about Training. UnixWorld’s Open Computing, 12(11):38–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Bruno:1995:PGI

[Bru95b] L. Bruno. The Proving Grounds: IBM and Hewlett Packard now boast sparkling technology migration centers where you can migrate your mainframe applications to open systems. Unix-World’s Open Computing, 12(8): 50–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Bruno:1995:RS

[Bru95c] Lee Bruno. A Ready-to-Use Server. Open Computing, 12(2): 87–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1995:ALN

[Bru95d] Lee Bruno. Administering Lotus Notes. Open Computing, 12(1): 73–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1995:RBP

[Bru95e] Lee Bruno. Reshaping Business Processes. Open Computing, 12 (1):97–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Bruno:1995:WWI

[Bru95f] Lee Bruno. When Wireless Is Worth It. Open Computing, 12(4): 57–??, April 1995. CODEN OP-COEB. ISSN 1078-2370.

Burgard:1991:NWL

[Bur91a] Mike Burgard. A New Way to License Software. UNIX/world, 8 (12):60–??, December 1991. ISSN 0739-5922.

Burgard:1991:MTT

[Bur91b] Mike Burgard. More than a Toy but No Bargain. UNIX/world, 8 (12):75–??, December 1991. ISSN 0739-5922.

Burgard:1991:RPT

[Bur91c] Mike Burgard. Not Ready for Prime Time. UNIX/world, 8(9): 105–??, September 1991. ISSN 0739-5922.

Burgard:1991:QD

[Bur91d] Mike Burgard. Quick on the Draw.

UNIX/world, 8(10):101–??, October 1991. ISSN 0739-5922.

Burgard:1991:SXP

[Bur91e] Mike Burgard. Serving Up X from a PC. UNIX/world, 8(11):99–??, November 1991. ISSN 0739-5922.

Burgard:1992:BSW

[Bur92a] Mike Burgard. Back-up Software Works While You Sleep.

UNIX/world, IX(12):103–??, December 1992. ISSN 0739-5922.

Burgard:1992:BYL

[Bur92b] Mike Burgard. Brought to You in Living Color. UNIX/world, 9(10): 77–??, October 1992. ISSN 0739-5922.

Burgard:1992:BTT

[Bur92c] Mike Burgard. Bull Tries the Turnkey Approach. UNIX/world, 9(2):67–??, February 1992. ISSN 0739-5922.

Burgard:1992:CU

[Bur92d] Mike Burgard. Counting by Unix. UNIX/world, IX(12):71–??, December 1992. ISSN 0739-5922.

Burgard:1992:DCS

[Bur92e] Mike Burgard. DECwrite comes to the SPARCstation. UNIX/world, 9(10):89–??, October 1992. ISSN 0739-5922.

Burgard:1992:LBY

[Bur92f] Mike Burgard. Look Before You Back Up. UNIX/world, 9(9):71– ??, September 1992. ISSN 0739-5922.

Burgard:1992:MMY

[Bur92g] Mike Burgard. Making the Most Out of Your Data. UNIX/world, 9 (9):85–??, September 1992. ISSN 0739-5922.

Burgard:1992:PM

[Bur92h] Mike Burgard. Product of the Month. UNIX/world, IX(12):41– ??, December 1992. ISSN 0739-

Burgard:1992:RS

[Bur92i] Mike Burgard. Real-Time Spreadsheets-Really! UNIX/world, 9(4):81–??, April 1992. ISSN 0739-5922.

Burgard:1992:SS

[Bur92j] Mike Burgard. Scanning for Scanners. UNIX/world, 9(7):72–??, July 1992. ISSN 0739-5922.

Burgard:1992:TCY

[Bur92k] Mike Burgard. Take Control of Your System. UNIX/world, 9(3): 75–??, March 1992. ISSN 0739-5922.

Burgard:1993:UEI

[Bur93a] Michael Burgard. UNIX E-mail is Growing Up. UNIX/world, 10(3): 111–??, March 1993. ISSN 0739-5922.

Burgard:1993:CEI

[Bur93b] Mike Burgard. Cross-Platform Email is a Reality. UNIX/world, 10 (1):101–??, January 1993. ISSN 0739-5922.

Burgard:1993:WFO

[Bur93c] Mike Burgard. When the Fax Overwhelms You. UNIX/world, 10(2):105–??, February 1993. ISSN 0739-5922.

Campbell:1991:EBU

[Cam91] Art Campbell. E-Mail Beyond UNIX. UNIX/world, 8(11):77–??, November 1991. ISSN 0739-5922.

Campbell:1992:MBP

[Cam92] Art Campbell. Mastering Business Presentations. UNIX/world, 9(1):

93–??, January 1992. ISSN 0739-5922.

Campbell:1995:USTa

[Cam95a] J. F. Campbell. Using small trucks to circumvent large truck restrictions: impacts on truck emissions and performance measures. Unix-World’s Open Computing, 12(12): 445–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Campbell:1995:USTb

[Cam95b] J. F. Campbell. Using small trucks to circumvent large truck restrictions: impacts on truck emissions and performance measures. Unix-World’s Open Computing, 12(12): 445–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Campbell:1995:USTc

[Cam95c] J. F. Campbell. Using small trucks to circumvent large truck restrictions: impacts on truck emissions and performance measures. Unix-World’s Open Computing, 12(12): 445–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Cardell:1995:DMC

[CHB95] N. S. Cardell, S. Huang, and S. Brown. Decisionmaking in the motor carrier industry. Unix-World’s Open Computing, 12(12): 401–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Clark:1994:EWU

[Cla94] David Clark. Easy Windows to Unix E-Mail. UnixWorld’s Open Computing, 11(9):74–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Cart-Mitchell:1992:BIF

[CMQ92] Smoot Cart-Mitchell and John S.

Quarterman. Building Internet Firewalls. UNIX/world, 9(2):93– ??, February 1992. ISSN 0739-5922.

Cortino:1994:YCK

[Cor94] Juli Cortino. You Can’t Kick the Tires of an RDBMS. Open Computing, 11(11):81–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Core:1995:BOB

[Cor95a] G. Core. Banking on Objects: The Bank of Nova Scotia depends on Versant ODBMS to help build its next generation of database applications that will support complex data types and increase worker productivity. UnixWorld’s Open Computing, 12(10):74–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Core:1995:LWH

[Cor95b] G. Core. Loans on the Web:

How Fannie Mae is using Web-Force, Silicon Graphics Inc’s Web authoring tool, to stake out a presence on the Internet. Unix-World’s Open Computing, 12(11): 74–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Crane:1994:SGF

[CR94] Elizabeth Crane and Richard Raucci. Say Goodbye to the File Cabinet. Open Computing, 11 (12):104–??, December 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Crane:1995:GD

[CR95] Elizabeth Crane and Richard Raucci. Going Digital. Open Computing, 12(2):93–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Crane:1994:BMS

[Cra94a] Elizabeth Crane. Back Me Up, Scotty. Open Computing, 11(12): 102–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Crane:1994:DCG

[Cra94b] Elizabeth E. Crane. The Data Conversion Gamble. Open Computing, 11(11):101–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Crane:1995:TPFb

[Cra95a] E. Crane. Telecommunications, Played Forte. UnixWorld’s Open Computing, 12(7):78–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Crane:1995:FSC

[Cra95b] Elizabeth Crane. Finding Some Comfort on the Web. Open Computing, 12(2):90–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Crane:1995:NR

[Cra95c] Elizabeth Crane. Notes in Residence. Open Computing, 12(1): 100–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Crane:1995:OWP

[Cra95d] Elizabeth Crane. OS/ 2 Warp Proves Its Point. Open Computing, 12(4):90–??, April 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Crane:1995:SNO

[Cra95e] Elizabeth Crane. Seeing New Opportunities. Open Computing, 12 (3):105–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Crane:1995:TPFa

[Cra95f] Elizabeth Crane. Telecommunications, Played Forte. Open Computing, 12(7):78–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Crane:1995:TSF

[Cra95g] Elizabeth Crane. Time/Sunsoft Firewall-1: How Safe is Safe? Open Computing, 12(5):92–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Diamond:1995:WTM

[Dia95] D. Diamond. Why Training Matters. UnixWorld’s Open Computing, 12(11):32–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Dodge:1994:RCa

[Dod94a] Marc Dodge. Reality Check. Open Computing, 11(10):29–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1994:RCb

[Dod94b] Marc Dodge. Reality Check. Open Computing, 11(11):31–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1994:RCc

[Dod94c] Marc Dodge. Reality Check. Open Computing, 11(12):27–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1994:RCd

[Dod94d] Marc Dodge. Reality Check. Unix-World’s Open Computing, 11(9): 31–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Dodge:1995:RCa

[Dod95a] Marc Dodge. Reality Check. Open Computing, 12(1):27–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1995:RCb

[Dod95b] Marc Dodge. Reality Check. Open Computing, 12(2):27–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1995:RCc

[Dod95c] Marc Dodge. Reality Check. Open Computing, 12(3):25–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1995:RCd

[Dod95d] Marc Dodge. Reality Check. Open Computing, 12(4):25–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1995:RCe

[Dod95e] Marc Dodge. Reality Check. Open Computing, 12(5):27–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1995:RCf

[Dod95f] Marc Dodge. Reality Check. Open Computing, 12(6):23–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Dodge:1995:RCg

[Dod95g] Marc Dodge. Reality Check.

Open Computing, 12(7):25–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Engler:1992:WCF

[Eng92] Natalie Engler. Wordperfect’s Change of Face. UNIX/world, 9 (11):87–??, November 1992. ISSN 0739-5922.

Engler:1993:NBL

[Eng93] Natalie Engler. No Business Like Show Business. UNIX/world, 10 (3):94–??, March 1993. ISSN 0739-5922.

Engler:1994:BEI

[Eng94a] Natalie Engler. Buying E-Mail Is Harder Than Ever. Unix-World’s Open Computing, 11(9): 88–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Engler:1994:DS

[Eng94b] Natalie Engler. Diane Smigel.

Open Computing, 11(12):56–??, December 1994. CODEN OP-COEB. ISSN 1078-2370.

Engler:1994:HIW

[Eng94c] Natalie Engler. Help is Where You Find It. Open Computing, 11(12): 62–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Engler:1994:DD

[Eng94d] Natalie Engler. The Dollars are in the Details. Open Computing, 11(10):42–??, October 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Engler:1995:DNC

[Eng95a] N. Engler. 3-D on the Desktop: As new capabilities bring 3-D down to workstations and PCs, managers of IS have one big question: Can it help my business? Unix-World’s Open Computing, 12(11): 42–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Engler:1995:IINb

[Eng95b] N. Engler. Is It Negligence? With the increased incidence of insiders using computers and networks to heist trade secrets and other intellectual property, many companies are still not going far enough to protect themselves. Unix-World’s Open Computing, 12(7): 44–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Engler:1995:IUS

[Eng95c] N. Engler. Is Uncle Sam Still Open?: Suddenly, the Fed’s commitment to open systems is becoming a marriage of convenience. UnixWorld’s Open Computing, 12 (10):36–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Engler:1995:LME

[Eng95d] N. Engler. Learning with Multimedia: Education Takes the Next Step. UnixWorld’s Open Computing, 12(11):36–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Engler:1995:PTO

[Eng95e] N. Engler. The Politics of Trust: Okay, so you’ve bought the groupware technology. Unix-World’s Open Computing, 12(9):

38–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Engler:1995:VES

[Eng95f] N. Engler. Videoconferencing:

Expect to see a boom in videoconferencing next year, as several models of PCs ship with full audio- and video-processing capabilities. UnixWorld’s Open Computing, 12(12):56–??, 1995. CO-DEN OPCOEB. ISSN 1072-4044.

Engler:1995:ATG

[Eng95g] Natalie Engler. Advanced Technology Groups: Set Up for Failure? Open Computing, 12(6): 38–??, June 1995. CODEN OP-COEB. ISSN 1078-2370.

Engler:1995:HSD

[Eng95h] Natalie Engler. How to Star in a Dysfunctional Company. Open Computing, 12(5):30–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Engler:1995:IINa

[Eng95i] Natalie Engler. Is It Negligence? Open Computing, 12(7): 44–??, July 1995. CODEN OP-COEB. ISSN 1078-2370.

Engler:1995:IMU

[Eng95j] Natalie Engler. IS Managers Under Stress. Open Computing, 12 (1):44–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Engler:1995:PS

[Eng95k] Natalie Engler. Power Shifting. Open Computing, 12(4):36– ??, April 1995. CODEN OP-COEB. ISSN 1078-2370.

Engler:1995:RBE

[Eng95l] Natalie Engler. Riding the Bleeding Edge of Distributed Computing. Open Computing, 12(2):34– ??, February 1995. CODEN OP-COEB. ISSN 1078-2370.

Engler:1995:RW

[Eng95m] Natalie Engler. The (Re)Movable Worker. Open Computing, 12(5): 40–??, May 1995. CODEN OP-COEB. ISSN 1078-2370.

Farris:1991:CMU

[Far91a] Rick Farris. Choosing a Modem for UNIX. UNIX/world, 8(10): 85–??, October 1991. ISSN 0739-5922.

Farris:1991:GYF

[Far91b] Rick Farris. Get Yer Free Software Here! UNIX/world, 8(12):95–??, December 1991. ISSN 0739-5922.

Farris:1991:MYC

[Far91c] Rick Farris. Must You Choose Between X and Windows? UNIX/world, 8(11):56–??, November 1991. ISSN 0739-5922.

Farris:1991:UMG

[Far91d] Rick Farris. UNIX Mail Gets Easier. UNIX/world, 8(11):111–??, November 1991. ISSN 0739-5922.

Farris:1991:WYR

[Far91e] Rick Farris. Would You Rather Have a Sun? UNIX/world, 8(10): 95–??, October 1991. ISSN 0739-5922.

Farrow:1991:WL

[Far91f] Rik Farrow. A Workstation for the Lap. UNIX/world, 8(9):97–??, September 1991. ISSN 0739-5922.

Farrow:1991:WSS

[Far91g] Rik Farrow. Will Success Spoil the Internet? UNIX/world, 8(9):79– ??, September 1991. ISSN 0739-5922.

Farris:1992:OLA

[Far92a] Rick Farris. An Open Look at 1-2-3. UNIX/world, 9(3):71–??, March 1992. ISSN 0739-5922.

Farris:1992:GCE

[Far92b] Rick Farris. Getting Connected with Ethernet. UNIX/world, 9(4): 65–??, April 1992. ISSN 0739-5922.

Farris:1992:MYP

[Far92c] Rick Farris. Maximizing Your PC for Graphics. UNIX/world, 9(2): 61–??, February 1992. ISSN 0739-5922.

Farris:1992:MEU

[Far92d] Rick Farris. Modern Editors for UNIX. UNIX/world, 9(2):73–??, February 1992. ISSN 0739-5922.

Farris:1992:RR

[Far92e] Rick Farris. RAID to the Rescue.

UNIX/world, 9(8):75–??, August 1992. ISSN 0739-5922.

Farris:1992:XRU

[Far92f] Rick Farris. X for the Rest of Us.

UNIX/world, 9(9):46–??, September 1992. ISSN 0739-5922.

Farrow:1992:AUS

[Far92g] Rik Farrow. Auditing UNIX Security, Effortlessly. UNIX/world, 9(8):85–??, August 1992. ISSN 0739-5922.

Farrow:1992:HIY

[Far92h] Rik Farrow. How to Improve Your System Security. UNIX/world, 9 (4):59–??, April 1992. ISSN 0739-5922.

Farrow:1992:LPS

[Far92i] Rik Farrow. Lower Price 3-D From Silicon Graphics. UNIX/world, 9(3):65–??, March 1992. ISSN 0739-5922.

Farrow:1992:NHU

[Far92j] Rik Farrow. New Hope for UNIX Security? UNIX/world, 9(10): 73–??, October 1992. ISSN 0739-5922.

Farrow:1992:ONM

[Far92k] Rik Farrow. Object-Oriented Network Management. UNIX/world, 9(11):93–??, November 1992. ISSN 0739-5922.

Farrow:1992:PV

[Far92l] Rik Farrow. Persistence of Vision. UNIX/world, IX(12):61–??, December 1992. ISSN 0739-5922.

Farrow:1992:TVS

[Far92m] Rik Farrow. Two Video Solutions for SPARCstations. UNIX/world, 9(4):69–??, April 1992. ISSN 0739-5922.

Farrow:1993:HW

[Far93a] Rik Farrow. Hooking Up to the World. UNIX/world, 10(3):61–??, March 1993. ISSN 0739-5922.

Farrow:1993:UWN

[Far93b] Rik Farrow. Understanding Windows NT. UNIX/world, 10(2):47– ??, February 1993. ISSN 0739-5922.

Farrow:1994:DTU

[Far94a] Rik Farrow. A Decision-Support Tool Under Control. Open Computing, 11(11):105–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Farrow:1994:SFT

[Far94b] Rik Farrow. A Switch for the Fast Track. Open Computing, 11(10): 87–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Farrow:1994:BAW

[Far94c] Rik Farrow. Barbarians at the Walls. Open Computing, 11(10): 62–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Farrow:1994:BMH

[Far94d] Rik Farrow. Breaking the Mainframe Habit. Open Computing, 11(10):93–??, October 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Farrow:1994:CUB

[Far94e] Rik Farrow. Cheap and Uneasy: BSD/ 386 and Linux. Unix-World’s Open Computing, 11(9): 83–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Farrow:1994:PYN

[Far94f] Rik Farrow. Protecting Your Network. Open Computing, 11(10): 83–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Farrow:1994:TV

[Far94g] Rik Farrow. Turn Up the Volume. Open Computing, 11(10):71– ??, October 1994. CODEN OP-COEB. ISSN 1078-2370.

Feldman:1995:NSN

[Fel95] Robert Feldman. Never Say Never to ATM. Open Computing, 12(3): 55–??, March 1995. CODEN OP-COEB. ISSN 1078-2370.

Farrow:1992:PUE

[FH92] Rik Farrow and Frank Hayes. The PC UNIX Explosion. UNIX/world, 9(10):48–??, October 1992. ISSN 0739-5922.

Fryer:1994:WYL

[Fry94] Bronwyn Fryer. “What Are You Ladies Doing Here?”. Open Computing, 11(12):34–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Fryer:1995:WOS

[Fry95a] B. Fryer. Will the Open Systems Crowd Embrace OLE? Unix-World’s Open Computing, 12(8): 54–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Fryer:1995:TCC

[Fry95b] Bronwyn Fryer. Three-Tier Client-Server Computing? Open Computing, 12(3):70–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Fryer:1994:OCT

[FS94] Bronwyn Fryer and Roderick Simpson. Open Computing’s Top 100 Women in Computing. Open Computing, 11(12):47–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Garner:1995:PQT

[Gar95a] R. Garner. Passing the Quality Test: A new study says clientserver apps have twice the defects of mainframe apps. Unix-World’s Open Computing, 12(10): 30–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Garner:1995:DMB

[Gar95b] Rochelle Garner. Do Management Books Help? Open Computing, 12(6):44–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Garner:1995:MM

[Gar95c] Rochelle Garner. Management Meltdown! Open Computing, 12 (1):36–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Garner:1995:MO

[Gar95d] Rochelle Garner. Managing Obsolescence. Open Computing, 12(3): 42–??, March 1995. CODEN OP-COEB. ISSN 1078-2370.

Garner:1995:GPM

[Gar95e] Rochelle Garner. The Growing Professional Menace. Open Computing, 12(7):32–??, July 1995. CODEN OPCOEB. ISSN 1078-

Garner:1995:TH

[Gar95f] Rochelle Garner. Turning Up the Heat. Open Computing, 12(4): 52–??, April 1995. CODEN OP-COEB. ISSN 1078-2370.

Gill:1995:SBB

[Gil95] P. J. Gill. Sear’s Big Budget Epic: Arbor Software’s Essbase multidimensional database helped Sears Roebuck and Co. consolidate its financial planning, analysis, and budgeting systems and repositories. UnixWorld’s Open Computing, 12(12):62–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Gomes:1992:SYV

[Gom92] Lee Gomes. Solving Your Vendor Problems Before They Occur. UNIX/world, IX(12):65–??, December 1992. ISSN 0739-5922.

Gould:1994:A

[Gou94a] Michael Gould. In Analysis. Unix-World’s Open Computing, 11(9): 37–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Goulde:1994:Aa

[Gou94b] Michael Goulde. In Analysis. Open Computing, 11(10):35–??, October 1994. CODEN OP-COEB. ISSN 1078-2370.

Goulde:1994:Ab

[Gou94c] Michael Goulde. In Analysis. Open Computing, 11(11):35–??, November 1994. CODEN OP-COEB. ISSN 1078-2370.

Goulde:1994:Ac

[Gou94d] Michael Goulde. In Analysis. Open Computing, 11(12):29–??, December 1994. CODEN OP-COEB. ISSN 1078-2370.

Goulde:1995:Aa

[Gou95a] Michael Goulde. In Analysis. Open Computing, 12(1):31– ??, January 1995. CODEN OP-COEB. ISSN 1078-2370.

Goulde:1995:Ab

[Gou95b] Michael Goulde. In Analysis. Open Computing, 12(2):29– ??, February 1995. CODEN OP-COEB. ISSN 1078-2370.

Goulde:1995:Ac

[Gou95c] Michael Goulde. In Analysis. Open Computing, 12(3):27–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Goulde:1995:Ad

[Gou95d] Michael Goulde. In Analysis. Open Computing, 12(4):27–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Goulde:1995:Ae

[Gou95e] Michael Goulde. In Analysis. Open Computing, 12(5):29– ??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Goulde:1995:Af

[Gou95f] Michael Goulde. In Analysis. Open Computing, 12(6):25–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Goulde:1995:Ag

[Gou95g] Michael Goulde. In Analysis. Open Computing, 12(7):27– ??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Granz:1992:SSE

[Gra92] David Granz. Storage Space in Epoch Proportions. UNIX/world, IX(12):77–??, December 1992. ISSN 0739-5922.

Granz:1993:DXA

[Gra93a] David Granz. Developing X Applications in DOS. UNIX/world, 10(2):99–??, February 1993. ISSN 0739-5922.

Granz:1993:WDW

[Gra93b] David Granz. Worldwide Dial-up with the Worldblazer. UNIX/world, 10(1):100–??, January 1993. ISSN 0739-5922.

Greenbaum:1992:FUC

[Gre92a] Joshua Greenbaum. France’s UNIX Champion. UNIX/world, 9 (9):61–??, September 1992. ISSN 0739-5922.

Greenbaum:1992:OOA

[Gre92b] Joshua Greenbaum. Open Oil at Elf Aquitaine. UNIX/world, 9 (10):55–??, October 1992. ISSN 0739-5922.

Greenbaum:1993:RUU

[Gre93]  Joshua Greenbaum.  Riding

UNIX into the U.S. Market. UNIX/world, 10(3):51–??, March 1993. ISSN 0739-5922.

Greenbaum:1994:MEa

[Gre94a] Joshua Greenbaum. Meanwhile, in Europe. Open Computing, 11 (10):37–??, October 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Greenbaum:1994:MEb

[Gre94b] Joshua Greenbaum. Meanwhile, in Europe. Open Computing, 11 (11):37–??, November 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Greenbaum:1994:MEc

[Gre94c] Joshua Greenbaum. Meanwhile, in Europe. Open Computing, 11 (12):31–??, December 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Greenbaum:1994:MEd

[Gre94d] Joshua Greenbaum. Meanwhile, in Europe. UnixWorld’s Open Computing, 11(9):39–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Greenbaum:1994:SVI

[Gre94e] Joshua Greenbaum. Silicon Valley, Israeli-Style. Open Computing, 11(10):50–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Greenbaum:1995:MEa

[Gre95a] Joshua Greenbaum. Meanwhile, in Europe. Open Computing, 12 (1):33–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Greenbaum:1995:MEb

[Gre95b] Joshua Greenbaum. Meanwhile, in Europe. Open Computing, 12 (2):31–??, February 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Greenbaum:1995:MEc

[Gre95c] Joshua Greenbaum. Meanwhile, in Europe. Open Computing, 12 (3):29–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Greenbaum:1995:MEd

[Gre95d] Joshua Greenbaum. Meanwhile, in Europe. Open Computing, 12 (4):29–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Greenbaum:1995:SAa

[Gre95e] Joshua Greenbaum. Service Agent. Open Computing, 12(6): 28–??, June 1995. CODEN OP-COEB. ISSN 1078-2370.

Greenbaum:1995:SAb

[Gre95f] Joshua Greenbaum. Service Agent. Open Computing, 12(7): 30–??, July 1995. CODEN OP-COEB. ISSN 1078-2370.

Guglielmo:1995:DS

[Gug95a] Connie Guglielmo. Documenting Success. Open Computing, 12(2): 57–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Guglielmo:1995:TF

[Gug95b] Connie Guglielmo. Taking Flight. Open Computing, 12(4):44–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Haber:1995:BPR

[Hab95] Lynn Haber. Business Process Reengineering Tools. Open Computing, 12(1):77–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Hayes:1991:GRW

[Hay91a] Frank Hayes. Getting Rid of Wires. UNIX/world, 8(10):70–??, October 1991. ISSN 0739-5922.

Hayes:1991:MCW

[Hay91b] Frank Hayes. Making CASE Work for You. UNIX/world, 8(11):70– ??, November 1991. ISSN 0739-5922.

Hayes:1991:UCB

[Hay91c] Frank Hayes. Unjamming Client-Server Bottlenecks. UNIX/world, 8(9):72–??, September 1991. ISSN 0739-5922.

Hayes:1992:EDM

[Hay92a] Frank Hayes. Exploding Desktop Myths. UNIX/world, 9(4):38–??, April 1992. ISSN 0739-5922.

Hayes:1992:FOS

[Hay92b] Frank Hayes. Facing the Open Systems Gap. UNIX/world, 9 (9):66–??, September 1992. ISSN 0739-5922.

Hayes:1992:GG

[Hay92c] Frank Hayes. Going Graphical. UNIX/world, 9(7):46–??, July 1992. ISSN 0739-5922.

Hayes:1992:OP

[Hay92d] Frank Hayes. Objective: Productivity. UNIX/world, 9(8):69–??, August 1992. ISSN 0739-5922.

Hayes:1992:PTU

[Hay92e] Frank Hayes. Programming Tools Unbundled. UNIX/world, 9(11): 65–??, November 1992. ISSN 0739-5922.

Hayes:1992:GD

[Hay92f] Frank Hayes. The Groupware Dilemma. UNIX/world, 9(2):46– ??, February 1992. ISSN 0739-

Hayes:1992:SAS

[Hay92g] Frank Hayes. The System Administration Squeeze. UNIX/world, 9 (10):67–??, October 1992. ISSN 0739-5922.

Hayes:1992:UCD

[Hay92h] Frank Hayes. UnixWorld’s Core Dump. UNIX/world, IX(12):58– ??, December 1992. ISSN 0739-5922.

Hayes:1993:ITL

[Hay93] Frank Hayes. Is There Life After NT? UNIX/world, 10(2):42–??, February 1993. ISSN 0739-5922.

Hayes:1993:MF

[HD93] Frank Hayes and Barbara Doheny. Market Forecast ’93. UNIX/world, 10(1):57–??, January 1993. ISSN 0739-5922.

Heath:1991:CAS

[Hea91] Robert Heath. Controlling Access to SVR4. UNIX/world, 8(11):123– ??, November 1991. ISSN 0739-5922.

Hein:1992:GYS

[Hei92] Trent R. Hein. Giving Your Sun the SLIP. UNIX/world, 9(4):91– ??, April 1992. ISSN 0739-5922.

Hirsch:1993:UOD

[Hir93] Phil Hirsch. Using OSF’s DCE. UNIX/world, 10(2):62–??, February 1993. ISSN 0739-5922.

Hayes:1993:NR

[HP93] Frank Hayes and Gary Andrew Poole. Novell to the Rescue.

UNIX/world, 10(3):42–??, March 1993. ISSN 0739-5922.

Kirzner:1995:PYC

[KB95] R. Kirzner and L. Bruno. 100 Products You Can’t Ignore in 1996: Looking for the jewels amid the computer industry’s characteristic smoke-and-mirrors? Unix-World’s Open Computing, 12(12): 27–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

King:1992:XIK

[Kin92] Peggy King. X.25 is the Key to Global WANs. UNIX/world, 9 (11):81–??, November 1992. ISSN 0739-5922.

King:1993:PS

[Kin93a] Peggy King. Product Snapshots.

UNIX/world, 10(1):111–??, January 1993. ISSN 0739-5922.

King:1993:WGC

[Kin93b] Peggy King. When Groupware Comes to UNIX. UNIX/world, 10(3):57–??, March 1993. ISSN 0739-5922.

Kirzner:1995:ETP

[Kir95a] R. Kirzner. External Training Packages: Not a Complete Solution. UnixWorld’s Open Computing, 12(11):35–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Kirzner:1995:HICb

[Kir95b] R. Kirzner. Help! I’m Caught in a Time Warp! Unix-World’s Open Computing, 12(7): 69–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Kirzner:1995:OFT

[Kir95c] R. Kirzner. Objects: Farther than They Appear. UnixWorld’s Open Computing, 12(11):55–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Kirzner:1995:PBIb

[Kir95d] R. Kirzner. Products to Block the Internal Threat. Unix-World’s Open Computing, 12(7): 48–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Kirzner:1995:PKBb

[Kir95e] R. Kirzner. Products to Keep the Barbarians O* U and T! Unix-World’s Open Computing, 12(7): 36–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Kirzner:1995:SYB

[Kir95f] R. Kirzner. Should You Bypass Windows 95? UnixWorld’s Open Computing, 12(10):57–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Kirzner:1995:CP

[Kir95g] Rikki Kirzner. Consuming Passions. Open Computing, 12(2):78– ??, February 1995. CODEN OP-COEB. ISSN 1078-2370.

Kirzner:1995:HICa

[Kir95h] Rikki Kirzner. Help! I’m Caught in a Time Warp! Open Computing, 12(7):69–??, July 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Kirzner:1995:IFT

[Kir95i] Rikki Kirzner. Internet: Freeway to Toll Road. Open Computing, 12(6):89–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Kirzner:1995:MWW

[Kir95j] Rikki Kirzner. Must We Wait for the 25th Century? Open Computing, 12(4):79–??, April 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Kirzner:1995:OS

[Kir95k] Rikki Kirzner. Open Season. Open Computing, 12(1):83–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Kirzner:1995:PBIa

[Kir95l] Rikki Kirzner. Products to Block the Internal Threat. Open Computing, 12(7):48–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Kirzner:1995:PKBa

[Kir95m] Rikki Kirzner. Products to Keep the Barbarians O*U and T! Open Computing, 12(7):36–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Kirzner:1995:HCD

[Kir95n] Rikki Kirzner. The High Cost of Dreaming. Open Computing, 12(5):75–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Klaus:1993:CL

[Kla93] Todd C. Klaus. Checking Out Linux. UNIX/world, 10(3):66–??, March 1993. ISSN 0739-5922.

Koland:1995:PDA

[Kol95a] Cordell Koland. Personal Digital Assistants? Open Computing, 12 (1):69–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Koland:1995:SSH

[Kol95b] Cordell Koland. Sharing or Segregating Host Resources. Open Computing, 12(2):65–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Krick:1995:RD

[Kri95] John Krick. Relating Documents. Open Computing, 12(1): 102–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Lacy:1992:LAG

[Lac92] Janice Ann Lacy. Leap Ahead in GUI Development. UNIX/world, 9(1):87–??, January 1992. ISSN 0739-5922.

LaPlante:1994:EO

[LaP94] Alice G. LaPlante. Everybody Objects. Open Computing, 11(11): 50–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Lash:1995:TCCb

[Las95a] A. Lash. Time to Call in the Cybertroops. UnixWorld’s Open Computing, 12(7):46–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Lash:1995:TCCa

[Las95b] Alex Lash. Time to Call in the Cybertroops. Open Computing, 12(7):46–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Leibowitz:1992:SWW

[Lei92] Michael R. Leibowitz. Still Waiting for the Windfall. UNIX/world, 9(2):53–??, February 1992. ISSN 0739-5922.

Leisner:1993:TUA

[Lei93] Marty Leisner. Tutorial: Using the AMD Automounter. UNIX/world, 10(1):70–??, January 1993. ISSN 0739-5922.

Lenatti:1995:RPM

[Len95a] C. Lenatti. Rethinking in Parallel: Multiprocessing is on the rise, despite a dearth of tools to help create multithreaded applications. UnixWorld’s Open Computing, 12(8):57–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Lenatti:1995:SMS

[Len95b] C. Lenatti. System Management Software: Want to manage PC desktops in a heterogeneous environment? Expect to spend the rest of your life in vendor meetings. UnixWorld’s Open Computing, 12(9):63–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Lenatti:1995:CVS

[Len95c] Chuck Lenatti. Commercial Vs.

Shareware TCP/IP Stacks. Open Computing, 12(6):54–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Lenatti:1995:FHB

[Len95d] Chuck Lenatti. Finding a Headhunter Before One Finds You. Open Computing, 12(5):46–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Linthicum:1994:KFP

[Lin94a] David S. Linthicum. Keeping a Flat Profile. Open Computing, 11 (11):67–??, November 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Linthicum:1994:IP

[Lin94b] David S. Linthicum. Out of Isolation, into Port. Open Computing, 11(12):84–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Linthicum:1994:TCY

[Lin94c] David S. Linthicum. Tommy, Can You See Me? UnixWorld’s Open Computing, 11(9):67–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Linthicum:1995:NMD

[Lin95a] D. S. Linthicum. Network Management in the Distributed Enterprise. UnixWorld’s Open Computing, 12(9):49–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Linthicum:1995:RFF

[Lin95b] D. S. Linthicum. RAD: Fast Food for Application Hunger. Unix-World’s Open Computing, 12(12): 48–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Linthicum:1995:TMT

[Lin95c] D. S. Linthicum. TP Monitors: Transaction processing monitors provide the management services that make client-server database computing trustworthy. Unix-World’s Open Computing, 12(7): 66–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Linthicum:1995:GDT

[Lin95d] David Linthicum. GUI Development Tools. Open Computing, 12 (3):73–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Linthicum:1995:WWF

[Lin95e] David Linthicum. Will Windows 95 Force a PC Upgrade? Open Computing, 12(4):76–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Linthicum:1995:CCD

[Lin95f] David S. Linthicum. Chasing Compound Document Standards. Open Computing, 12(1):55– ??, January 1995. CODEN OP-COEB. ISSN 1078-2370.

Linthicum:1995:C

[Lin95g] David S. Linthicum. CORBA 2.0? Open Computing, 12(2):68– ??, February 1995. CODEN OP-COEB. ISSN 1078-2370.

Linthicum:1995:TM

[Lin95h] David S. Linthicum. TP Monitors. Open Computing, 12(7):66– ??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Maguire:1995:HJH

[Mag95a] J. G. Maguire. Hot Java for Homer Simpson: Sun Microsystems Inc revives its interactive technology for Fox Broadcasting Co’s 3-D Web site. UnixWorld’s Open Computing, 12(11):72–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Maguire:1995:NWN

[Mag95b] J. G. Maguire. No Wrong Numbers with Objectivity: With the help of Objectivity/DB, Motorola is building a global, satellite-based cellular telephone network that will bring telephone service into areas where no telephone company has previously ventured. Unix-World’s Open Computing, 12(10): 72–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Maguire:1995:MCT

[Mag95c] John G. Maguire. McCaw Cellular Takes the NextStep. Open Computing, 12(4):93–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Maguire:1995:MCN

[Mag95d] John G. Maguire. Mercury Center/Netscape’s Netsite Server: Attracting Internet Customers. Open Computing, 12(5):94–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Mansisidor:1995:VP

[Man95] Julie Mansisidor. Visionary Programming. Open Computing, 12 (3):99–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Moore:1992:FCC

[Moo92a] Tom Moore. Faster Color Comes to the Black Next. UNIX/world, 9 (9):77–??, September 1992. ISSN 0739-5922.

Moore:1992:LWF

[Moo92b] Tom Moore. Low-end Workstations with Fangs. UNIX/world, 9(8):81–??, August 1992. ISSN 0739-5922.

Mikes:1991:FD

[MS91] Steven Mikes and Alan Southerton. Friendly Desktops. UNIX/world, 8(9):52–??, September 1991. ISSN 0739-5922.

Mui:1992:IXW

[Mui92] Linda Mui. Improving X Windows Security. UNIX/world, IX(12): 115–??, December 1992. ISSN 0739-5922.

Mullich:1995:SMB

[Mul95a] J. Mullich. Squashing the Millennium Bug: Think the Millenium Bug problem is something you can ignore? UnixWorld’s Open Computing, 12(10):42–??, 1995. CO-DEN OPCOEB. ISSN 1072-4044.

Mullich:1995:RMI

[Mul95b] J. Mullich. The Riddle in the Middle: It’s expensive, it’s complicated, and its implementation is a barometer of a company’s aversion to risk. UnixWorld’s Open Computing, 12(8):36–??, 1995. CO-DEN OPCOEB. ISSN 1072-4044.

Mullich:1995:FCT

[Mul95c] Joe Mullich. Fat Clients, Thin Patience. Open Computing, 12(3): 50–??, March 1995. CODEN OP-COEB. ISSN 1078-2370.

Mullich:1995:INA

[Mul95d] Joe Mullich. It’s Nothing Automatic. Open Computing, 12(2): 46–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Murray:1991:HDA

[Mur91] Jim Murray. HP’s Dual Approach to Workstations. UNIX/world, 8 (11):91–??, November 1991. ISSN 0739-5922.

Mynatt:1992:TYW

[Myn92] Beth Mynatt. Translating Your Way Into GUIs. UNIX/world, 9 (4):75–??, April 1992. ISSN 0739-5922.

Nee:1991:GYP

[Nee91a] Eric Nee. Get Your PC Software on UNIX. UNIX/world, 8(10): 54–??, October 1991. ISSN 0739-5922.

Nee:1991:LSL

[Nee91b] Eric Nee. Life in the Slow Lane. UNIX/world, 8(10):60–??, October 1991. ISSN 0739-5922.

Nee:1992:DT

[Nee92] Eric Nee. DEC in Transition. UNIX/world, 9(4):47–??, April 1992. ISSN 0739-5922.

OConnor:1994:NN

[O’C94] Rory J. O’Connor. The Next Noorda? Open Computing, 11(10): 59–??, October 1994. CODEN OPCOEB. ISSN 1078-2370.

Perkins:1992:PIY

[Per92a] Edwin C. Perkins. The PC Inside Your Sun Workstation. UNIX/world, 9(10):83–??, October 1992. ISSN 0739-5922.

Perkins:1992:XWR

[Per92b] Edwin C. Perkins, Jr. X and Windows: Roommates of Your PC. UNIX/world, IX(12):97–??, December 1992. ISSN 0739-5922.

Perkins:1993:OMW

[Per93a] Edwin C. Perkins, Jr. Opening More Windows. UNIX/world, 10 (3):109–??, March 1993. ISSN 0739-5922.

Perkins:1993:SYN

[Per93b] Edwin C. Perkins, Jr. Should Your Next UNIX Workstation be a PC? UNIX/world, 10(2):90–??, February 1993. ISSN 0739-5922.

Pfister:1994:OCM

[Pfi94] Sean Pfister. Open Computing MIS Survey: Benchmarks Miss Their Mark. UnixWorld’s Open Computing, 11(9):50–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Poole:1991:IDG

[Poo91a] Gary Andrew Poole. Is Data General Headed for Another Crash? UNIX/world, 8(11):63–??, November 1991. ISSN 0739-5922.

Poole:1991:STE

[Poo91b] Gary Andrew Poole. Sun in Their Eyes. UNIX/world, 8(10):79–??, October 1991. ISSN 0739-5922.

Poole:1991:MN

[Poo91c] Gary Andrew Poole. The Mellowing of Next. UNIX/world, 8(9):65– ??, September 1991. ISSN 0739-5922.

Poole:1991:TUC

[Poo91d] Gary Andrew Poole. The Top 10 UNIX Companies of 1991. UNIX/world, 8(12):51–??, December 1991. ISSN 0739-5922.

Poole:1992:AHA

[Poo92a] Gary Andrew Poole. Andy Heller Against the World. UNIX/world, 9(3):34–??, March 1992. ISSN 0739-5922.

Poole:1992:BP

[Poo92b] Gary Andrew Poole. From Blue to Pink. UNIX/world, 9(10):61–??, October 1992. ISSN 0739-5922.

Poole:1992:MFO

[Poo92c] Gary Andrew Poole. Market Forecast: Open Systems in 1992. UNIX/world, 9(1):72–??, January 1992. ISSN 0739-5922.

Poole:1992:SSC

[Poo92d] Gary Andrew Poole. Sequent’s Software Cure. UNIX/world, 9(3): 41–??, March 1992. ISSN 0739-5922.

Poole:1992:SB

[Poo92e] Gary Andrew Poole. Sun Block. UNIX/world, 9(7):55–??, July 1992. ISSN 0739-5922.

Poole:1992:SSS

[Poo92f] Gary Andrew Poole. Sun’s Soft Spot. UNIX/world, 9(11):52–??, November 1992. ISSN 0739-5922.

Poole:1992:SSJ

[Poo92g] Gary Andrew Poole. Sybase Stays a Jump Ahead. UNIX/world, 9 (9):55–??, September 1992. ISSN 0739-5922.

Poole:1992:DPA

[Poo92h] Gary Andrew Poole. The Dealer Plays an ACE. UNIX/world, 9 (4):53–??, April 1992. ISSN 0739-5922.

Poole:1992:GC

[Poo92i] Gary Andrew Poole. The Great Conciliator. UNIX/world, 9(8): 59–??, August 1992. ISSN 0739-

Poole:1993:NHN

[Poo93a] Gary Andrew Poole. And Now … Heeeeere’s NT! UNIX/world, 10 (2):53–??, February 1993. ISSN 0739-5922.

Poole:1993:PBB

[Poo93b] Gary Andrew Poole. Profile: Big Blue’s Common Man. UNIX/world, 10(1):63–??, January 1993. ISSN 0739-5922.

Poole:1993:BBN

[Poo93c] Gary Andrew Poole. The Brain Behind NT. UNIX/world, 10(2): 59–??, February 1993. ISSN 0739-5922.

Poole:1994:HGT

[Poo94] Gary Andrew Poole. How to Get to the Top in IS. Open Computing, 11(11):42–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Poole:1995:AWT

[Poo95a] G. A. Poole. Avoiding Web Tofu:

You probably want to know what not to do to develop a corporate identity on the World Wide Web. UnixWorld’s Open Computing, 12 (9):53–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Poole:1995:LMM

[Poo95b] G. A. Poole. Logout: Mellowing machines. UnixWorld’s Open Computing, 12(12):80–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Poole:1995:TPC

[Poo95c] G. A. Poole. Training Partnerships: Can You Make Them Last?

UnixWorld’s Open Computing, 12 (11):39–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Poole:1995:La

[Poo95d] Gary Andrew Poole. Logout. Open Computing, 12(1):112–??, January 1995. CODEN OP-COEB. ISSN 1078-2370.

Poole:1995:Lb

[Poo95e] Gary Andrew Poole. Logout. Open Computing, 12(2):104–??, February 1995. CODEN OP-COEB. ISSN 1078-2370.

Poole:1995:Lc

[Poo95f] Gary Andrew Poole. Logout. Open Computing, 12(3):112–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Poole:1995:Ld

[Poo95g] Gary Andrew Poole. Logout. Open Computing, 12(4):104–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Poole:1995:Le

[Poo95h] Gary Andrew Poole. Logout. Open Computing, 12(5):104–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Poole:1995:Lf

[Poo95i] Gary Andrew Poole. Logout. Open Computing, 12(6):120–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Poole:1995:Lg

[Poo95j] Gary Andrew Poole. Logout. Open Computing, 12(7):96–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Pooler:1995:USS

[Poo95k] J. A. Pooler. The use of spatial separation in the measurement of transportation accessibility. Unix-World’s Open Computing, 12(12): 421–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Poole:1992:GUU

[PW92] Gary Andrew Poole and Carolyn W. C. Wong. Getting Use Out of User Groups. UNIX/world, 9(3): 55–??, March 1992. ISSN 0739-5922.

Ranum:1994:IFP

[Ran94] Marcus Ranum. Internet Firewall Protection. UnixWorld’s Open Computing, 11(9):95–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Ranade:1995:WMS

[Ran95] Jay Ranade. Why the Mainframe Survives. Open Computing, 12(4): 34–??, April 1995. CODEN OP-COEB. ISSN 1078-2370.

Raucci:1994:ORa

[Rau94a] Richard Raucci. On-Site Reviews. Open Computing, 11(11): 99–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Raucci:1994:ORb

[Rau94b] Richard Raucci. On-Site Reviews. Open Computing, 11(12): 99–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Raucci:1994:LL

[Rau94c] Richard Raucci. The Lexus of Laptops. UnixWorld’s Open Computing, 11(9):79–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Raucci:1995:MHCb

[Rau95a] R. Raucci. Modern Health Care. UnixWorld’s Open Computing, 12 (7):83–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Raucci:1995:MHCa

[Rau95b] Richard Raucci. Modern Health Care. Open Computing, 12(7):83– ??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Raucci:1995:OCN

[Rau95c] Richard Raucci. Old Cobol, New Techniques. Open Computing, 12 (6):107–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Raucci:1995:RS

[Rau95d] Richard Raucci. RAD in the Stars. Open Computing, 12(6): 110–??, June 1995. CODEN OP-COEB. ISSN 1078-2370. Reynolds-Feighan:1995:EAA

[RF95] A. J. Reynolds-Feighan. European and American approaches to air transport liberalisation: some implications for small communities. UnixWorld’s Open Computing, 12 (12):467–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Raney:1991:NAO

[RM91] Scott Raney and Elizabeth Mynatt. New Age Office Automation. UNIX/world, 8(12):81–??, December 1991. ISSN 0739-5922.

Roberts:1995:MUL

[Rob95] B. Roberts. Managing Upheaval: Like many information services executives, National Semiconductor Corp.’s Connie Deletis is orchestrating radical change. Unix-World’s Open Computing, 12(12): 42–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Rohm:1995:SAB

[Roh95] W. G. Rohm. Slow Adoption, Big Payoff: For Commerce Clearing House, Lotus Notes was more than a business solution. Unix-World’s Open Computing, 12(9): 47–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Ryan:1994:MMC

[Rya94] Pat Ryan. Making Mosaic Clients Talk Back. Open Computing, 11 (10):99–??, October 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Sandlund:1993:PS

[San93] Chris Sandlund. Product Snapshots. UNIX/world, 10(2):111–??, February 1993. ISSN 0739-5922.

Schlosberg:1995:OIC

[Sch95a] J. Schlosberg. The Object Is Complexity: Text-only relational databases are slowly giving way to a new class of object-oriented, or object relational, databases. Unix-World’s Open Computing, 12(10): 52–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Schlosberg:1995:WIE

[Sch95b] J. Schlosberg. Where is the Elusive Groupware Payoff? Unix-World’s Open Computing, 12(9): 30–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Southerton:1993:BP

[SE93] Alan Southerton and Natalie Engler. The Best Products of 1992. UNIX/world, 10(1):46–??, January 1993. ISSN 0739-5922.

Sheldrick:1992:SXW

[She92] Dennis Sheldrick. Security and the X Window System. UNIX/world, 9(1):103–??, January 1992. ISSN 0739-5922.

Simpson:1995:SBS

[Sim95a] Rod Simpson. Suites or Best-of-Breed Software. Open Computing, 12(3):61–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Simpson:1995:WSR

[Sim95b] Roderick Simpson. Why Salaries Are on the Rise. Open Computing, 12(5):53–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Snell:1992:IBO

[Sne92] Ned Snell. IBM Builds an Open Strategy. UNIX/world, 9(1):65– ??, January 1992. ISSN 0739-5922.

Spoutherton:1992:STP

[SP92] Alan Spoutherton and Edwin Perkins. Sun Tries to Play Catchup. UNIX/world, 9(11):57–??, November 1992. ISSN 0739-5922.

Stevens:1995:CRG

[Ste95a] L. Stevens. Civil Rights Group Frames Its Web Site. Unix-World’s Open Computing, 12(11): 68–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Stevens:1995:HMS

[Ste95b] L. Stevens. High-End Multiprocessing Servers. UnixWorld’s Open Computing, 12(8):67–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Stevens:1995:PSO

[Ste95c] L. Stevens. Polishing Sales with OLAP: Using the DSS Agent relational OLAP tool, S.C. Johnson Wax creates data reports to win more shelf space for its products. UnixWorld’s Open Computing, 12 (12):71–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Stevens:1995:USP

[Ste95d] L. Stevens. UniSQL Solves the Paper Chase. UnixWorld’s Open Computing, 12(10):68–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Stevens:1995:HSM

[Ste95e] Larry Stevens. Hierarchical Storage Management. Open Computing, 12(5):71–??, May 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Stevens:1995:MYB

[Ste95f] Larry Stevens. Mapping Your Business Processes. Open Computing, 12(2):73–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Stevens:1995:XF

[Ste95g] Larry Stevens. The X Factor. Open Computing, 12(6):61– ??, June 1995. CODEN OP-COEB. ISSN 1078-2370.

Streeter:1995:CCN

[Str95a] A. Streeter. Call Centers: Now that automatic call distributors link to many standard PBXs and LANs — and include new features — it may be time to invest in your call center. Unix-World’s Open Computing, 12(12): 52–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Streeter:1995:ISWb

[Str95b] A. Streeter. Is SATAN Worth It? UnixWorld’s Open Computing, 12(7):34–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Streeter:1995:HMT

[Str95c] April Streeter. Hey Mac, Take a Letter. Open Computing, 12(5): 55–??, May 1995. CODEN OP-COEB. ISSN 1078-2370.

Streeter:1995:ISWa

[Str95d] April Streeter. Is SATAN Worth It? Open Computing, 12(7):34– ??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Stump:1992:DM

[Stu92] Kean Stump. Delivering the Mail. UNIX/world, 9(11):103–??, November 1992. ISSN 0739-5922.

Swartz:1992:IGG

[Swa92] Ray Swartz. Island Graphics Grapples Along. UNIX/world, 9 (2):81–??, February 1992. ISSN 0739-5922.

Southerton:1992:UTS

[SWG92] Alan Southerton, Andrew D. Wolfe Jr., and David Granz. DOS and UNIX On a Two-Way Street. UNIX/world, 9(8):48–??, August 1992. ISSN 0739-5922.

Taylor:1993:WCS

[Tay93] Allen G. Taylor. What Are Customers Saying about Alpha? UNIX/world, 10(1):67–??, January 1993. ISSN 0739-5922.

Tristram:1994:BLM

[Tri94a] Claire Tristram. Buying or Leasing Midrange Servers. Open Computing, 11(11):71–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1994:GC

[Tri94b] Claire Tristram. Green Computers? Open Computing, 11(11):77– ??, November 1994. CODEN OP-COEB. ISSN 1078-2370.

Tristram:1994:MCA

[Tri94c] Claire Tristram. Migrating to Client-Server Accounting. Open Computing, 11(11):75–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1994:NRU

[Tri94d] Claire Tristram. Novell Rediscovers Unix. Open Computing, 11 (10):54–??, October 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Tristram:1994:PD

[Tri94e] Claire Tristram. PCMCIA on the Desktop. Open Computing, 11 (12):81–??, December 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Tristram:1994:WVW

[Tri94f] Claire Tristram. Windows vs.

Windows Emulation on Unix.

Open Computing, 11(12):72–??, December 1994. CODEN OP-COEB. ISSN 1078-2370.

Tristram:1994:VCS

[Tri94g] Clarie Tristram. Version Control Software? Open Computing, 11 (12):75–??, December 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Tristram:1995:BBS

[Tri95a] C. Tristram. Becoming a Beta Site or Not: It’s often a risky venture to join a beta program. Unix-World’s Open Computing, 12(9): 58–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:FCW

[Tri95b] C. Tristram. Forming a Committee… or Not: Who says groups always make better decisions than individuals? UnixWorld’s Open Computing, 12(12):54–??, 1995. CO-DEN OPCOEB. ISSN 1072-4044.

Tristram:1995:HLb

[Tri95c] C. Tristram. High-Speed LANs. UnixWorld’s Open Computing, 12 (7):59–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Tristram:1995:LMSb

[Tri95d] C. Tristram. License Management Software? UnixWorld’s Open Computing, 12(7):65–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:LSL

[Tri95e] C. Tristram. Load-Balancing Software: Load-balancing software promises to give you the performance clout of a supercomputer from the workstations you already have. UnixWorld’s Open Computing, 12(8):60–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:MSb

[Tri95f] C. Tristram. Microsoft SMS?

UnixWorld’s Open Computing, 12 (7):62–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Tristram:1995:SYR

[Tri95g] C. Tristram. Should You Retrain Programmers? UnixWorld’s Open Computing, 12(11):33–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:SYSc

[Tri95h] C. Tristram. Should You Standardize on Windows 95? Unix-World’s Open Computing, 12(10): 44–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:SVR

[Tri95i] C. Tristram. Switching Vs. Routing: Switches aren’t ready to replace routers yet-but in a few years you may not be able to tell the difference between the two. UnixWorld’s Open Computing, 12 (8):64–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Tristram:1995:TIL

[Tri95j] C. Tristram. TCP/IP on LANs:

You may think that standardizing with a single transport protocol will lower your support costs. UnixWorld’s Open Computing, 12 (11):46–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Tristram:1995:TLE

[Tri95k] C. Tristram. TSAPI?: To a large extent, it depends on whether or not you already have the right equipment installed-and which vendor you ask. Unix-World’s Open Computing, 12(12): 50–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:WMW

[Tri95l] C. Tristram. WOSA: Microsoft’s Windows Open Systems Architecture is supposed to be the tie that binds your enterprise. Unix-World’s Open Computing, 12(11): 48–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:DYRa

[Tri95m] Claire Tristram. Do You Really Need ISO 9000? Open Computing, 12(5):65–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:DYRb

[Tri95n] Claire Tristram. Do You Really (Still) Need … Cobol? Open Computing, 12(6):57–??, June 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:F

[Tri95o] Claire Tristram. Firewalls. Open Computing, 12(5):61–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:HLa

[Tri95p] Claire Tristram. High-Speed LANs. Open Computing, 12(7): 59–??, July 1995. CODEN OP-COEB. ISSN 1078-2370.

Tristram:1995:LMSa

[Tri95q] Claire Tristram. License Management Software? Open Computing, 12(7):65–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:MSa

[Tri95r] Claire Tristram. Microsoft SMS?

Open Computing, 12(7):62–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:NPM

[Tri95s] Claire Tristram. Network Performance Monitoring. Open Computing, 12(6):51–??, June 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Tristram:1995:OD

[Tri95t] Claire Tristram. Object Databases. Open Computing, 12(4): 72–??, April 1995. CODEN OP-COEB. ISSN 1078-2370.

Tristram:1995:SYSa

[Tri95u] Claire Tristram. Should You Standardize On … Netware 4.1? Open Computing, 12(4):63–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:SYSb

[Tri95v] Claire Tristram. Should You Standardize On … Simple Network Management Protocol? Open Computing, 12(5):68–??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:SU

[Tri95w] Claire Tristram. Spec 1170-based Unix. Open Computing, 12(3):65– ??, March 1995. CODEN OP-COEB. ISSN 1078-2370.

Tristram:1995:SMO

[Tri95x] Claire Tristram. Stalking the Mega Outsourcing Deal. Open Computing, 12(3):32–??, March 1995. CODEN OPCOEB. ISSN 1078-2370.

Tristram:1995:TTK

[Tri95y] Claire Tristram. Ten Things To Know About … Implementing ISDN. Open Computing, 12(4): 67–??, April 1995. CODEN OP-COEB. ISSN 1078-2370.

Tristram:1995:TTI

[Tri95z] Claire Tristram. The Trouble with TCP/IP. Open Computing, 12(2): 52–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Tweed:1994:CR

[Twe94] Vera Tweed. Carol Robbins. Open Computing, 11(12):53–??, December 1994. CODEN OPCOEB. ISSN 1078-2370.

Tweed:1995:SLC

[Twe95] Vera Tweed. Small or Large Color Printers. Open Computing, 12(1): 63–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Verhoef:1995:DCU

[Ver95] E. Verhoef. The demand curve under road pricing and the problem of political feasibility. Unix-World’s Open Computing, 12(12): 459–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Wall:1991:MPB

[Wal91a] Larry Wall. The Mystery of the Perl Bug. UNIX/world, 8(10):

111–??, October 1991. ISSN 0739-5922.

Wall:1991:PLR

[Wal91b] Larry Wall. The Perl of Least Resistance. UNIX/world, 8(9):113– ??, September 1991. ISSN 0739-5922.

Ward:1992:AYD

[War92] Christopher Ward. Archie, Your Directory for Internet Software. UNIX/world, 9(9):99–??, September 1992. ISSN 0739-5922.

Weston:1994:SSS

[Wes94a] Rusty Weston. He Said, She Said, They Said. Open Computing, 11 (12):42–??, December 1994. CO-DEN OPCOEB. ISSN 1078-2370.

Weston:1994:YRW

[Wes94b] Rusty Weston. Your Results Will Vary. UnixWorld’s Open Computing, 11(9):42–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Weston:1995:PLR

[Wes95] Rusty Weston. Pushing the Limits of RAD. Open Computing, 12(6): 30–??, June 1995. CODEN OP-COEB. ISSN 1078-2370.

Wolfe:1992:PM

[Wol92] Andrew D. Wolfe, Jr. Product of the Month. UNIX/world, 9 (11):47–??, November 1992. ISSN 0739-5922.

Wolfe:1993:TBP

[Wol93] Andrew D. Wolfe, Jr. Transform Brings Popular Data Visualization to UNIX. UNIX/world, 10(1):

107–??, January 1993. ISSN 0739-5922.

Wolf:1995:EIS

[Wol95] Gary Wolf. Establishing an Internet Storefront. Open Computing, 12(1):50–??, January 1995. CO-DEN OPCOEB. ISSN 1078-2370.

Wong:1992:TUC

[Won92] Carolyn W. C. Wong. The Top 10 UNIX Companies of 1992. UNIX/world, IX(12):46–??, December 1992. ISSN 0739-5922.

Wong:1993:PS

[Won93] Carolyn W. C. Wong. Product Snapshot. UNIX/world, 10(3): 117–??, March 1993. ISSN 0739-5922.

Wong:1994:DL

[Won94a] Carolyn W. C. Wong. Dawn Lepore. Open Computing, 11(12):51– ??, December 1994. CODEN OP-COEB. ISSN 1078-2370.

Wong:1994:WCS

[Won94b] Carolyn W. C. Wong. Will CBT Solve Your Training Woes? Open Computing, 11(11):58–??, November 1994. CODEN OPCOEB. ISSN 1078-2370.

Wong:1995:TSH

[Won95a] C. W. C. Wong. A Training Success: How Clark County, Nevada Did It. UnixWorld’s Open Computing, 12(11):37–??, 1995. CO-DEN OPCOEB. ISSN 1072-4044.

Wong:1995:CIW

[Won95b] C. W. C. Wong. CDE 1.0?: It’s a well-regarded Unix desktop.

Maybe it’s about seven years too late to matter. UnixWorld’s Open Computing, 12(8):62–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Wong:1995:VMU

[Won95c] C. W. C. Wong. Vendors Are from Mars, Users Are from Venus. UnixWorld’s Open Computing, 12 (8):44–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Wong:1995:WPL

[Won95d] C. W. C. Wong. What PG and E Learned about Training. Unix-World’s Open Computing, 12(11): 34–??, 1995. CODEN OPCOEB. ISSN 1072-4044.

Wong:1995:NPA

[Won95e] Carolyn W. C. Wong. Negotiating Perks for Advanced Dummies. Open Computing, 12(5):50– ??, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Wolfe:1993:NMX

[WP93] Andrew D. Wolfe, Jr. and Edwin C. Perkins, Jr. Next Meets X. UNIX/world, 10(3):101–??, March 1993. ISSN 0739-5922.

Yager:1994:HLL

[Yag94] Tom Yager. How I Learned to Live With Benchmarks. Unix-World’s Open Computing, 11(9): 56–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Zintz:1992:YWL

[Zin92] Walter Zintz. So, You Wanna Learn About UNIX. UNIX/world, 9(7):65–??, July 1992. ISSN 0739-

Zintz:1994:MIa

[Zin94a] Walter Zintz. Meet the Internet. Open Computing, 11(10):74– ??, October 1994. CODEN OP-COEB. ISSN 1078-2370.

Zintz:1994:MIb

[Zin94b] Walter Zintz. Meet the Internet. Open Computing, 11(11):113– ??, November 1994. CODEN OP-COEB. ISSN 1078-2370.

Zintz:1994:SPN

[Zin94c] Walter Zintz. Shrewd Projects; Nifty Utilities. UnixWorld’s Open Computing, 11(9):76–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Zintz:1995:La

[Zin95a] Walter Zintz. Library. Open Computing, 12(1):104–??, January 1995. CODEN OPCOEB. ISSN 1078-2370.

Zintz:1995:Lb

[Zin95b] Walter Zintz. Library. Open Computing, 12(2):96–??, February 1995. CODEN OPCOEB. ISSN 1078-2370.

Zintz:1995:Lc

[Zin95c] Walter Zintz. Library. Open Computing, 12(4):98–??, April 1995. CODEN OPCOEB. ISSN 1078-2370.

Zintz:1995:Ld

[Zin95d] Walter Zintz. Library. Open Computing, 12(7):87–??, July 1995. CODEN OPCOEB. ISSN 1078-2370.